### v1

# boxthm]Proposition boxthm]Example boxthm]Corollary boxthm]Lemma boxthm]Definition boxthm]Assumption

Abstract
Ensuring fairness in matching algorithms is a key challenge in allocating scarce resources and positions. Focusing on Optimal Transport (OT), we introduce a novel notion of group fairness requiring that the probability of matching two individuals from any two given groups in the OT plan satisfies a predefined target. We first propose FairSinkhorn, a modified Sinkhorn algorithm to compute perfectly fair transport plans efficiently. Since exact fairness can significantly degrade matching quality in practice, we then develop two relaxation strategies. The first one involves solving a penalised OT problem, for which we derive novel finite-sample complexity guarantees. This result is of independent interest as it can be generalized to arbitrary convex penalties. Our second strategy leverages bilevel optimization to learn a ground cost that induces a fair OT solution, and we establish a bound guaranteeing that the learned cost yields fair matchings on unseen data. Finally, we present empirical results that illustrate the trade-offs between fairness and performance.
1 Introduction
Algorithmic matching mechanisms play an increasing role in modern societies, handling the distribution of rare goods by connecting individuals or firms through algorithmic decisions rather than through price-driven markets. Instead of choosing an exchange partner through a decentralized market in which prices make matches, participants are allocated one or multiple partners through a centralized clearing house that takes into account their individual characteristics and preferences. Examples of such mechanisms include online job recommendations, college admissions systems, dating and ride-hailing apps, and kidney allocation circuits.
An increasing concern is the fairness of such mechanisms: since matchings partially determine individual outcomes (for instance, through marriage or schooling decisions), they may result in unfair outcomes. For instance, the growing importance of dating apps in contemporary matrimonial dynamics has raised concerns over their role in fostering social and racial marital homogamy [zheng2018fairness, jia2018online, zhao2024leveraging], a phenomenon often highlighted as a key driver of inequality. Similarly, school admission systems have been criticized for excessively matching students from privileged backgrounds with elite institutions regardless of their academic potential [hiller2012choix, fack2014impact, simioni2022comment]—causing underprivileged students to lose the benefits of elite higher education. More generally,
these concerns arise because matching decisions overly depend on individual characteristics that are highly correlated with attributes defining social groups, such as ethnicity or socioeconomic background.
Such dependence may lead to highly homogeneous pairings, which may conflict with political, philosophical, and legal principles that promote diversity and equality of opportunity. Algorithmic decisions systematically disadvantaging certain groups are often referred to as group fairness issues.
Group fairness has been extensively studied in supervised and unsupervised learning [barocas-hardt-narayanan]. In supervised learning, where an algorithm predicts an unknown outcome $Y$ from features $X$, one example of group fairness is demographic parity [calders2009building], which requires that the prediction $f(X)$ be independent of a protected attribute $S$ (such as gender or age). In practice, this means that predictions for individuals—such as credit scores—should not systematically depend on sensitive attributes. In unsupervised learning, one approach is equality of representation error, which requires that the reconstruction error of a learned latent representation be independent of the sensitive attribute $S$ [samadi2018price]. Other notions of group fairness have also been proposed in both supervised and unsupervised contexts [barocas-hardt-narayanan].
Extending these fairness notions to matching problems is, however, nontrivial, as matching cannot be directly framed as either a supervised or unsupervised learning task. Most prior work has focused on individual fairness (ensuring that similar individuals receive similar matches) or on problem-specific notions, such as guaranteeing a minimal participation level in settings where individuals can opt out of the matching mechanism. We discuss this literature in more details in the related work section.
Contributions.
In this work, we study group fairness in matching using the framework of optimal transport (OT). OT has been a long standing modeling tool in economics and social sciences since the pioneering work of Kantorovitch, eventually earning him the Nobel Memorial Prize in Economic Sciences in 1975 [galichon2018optimal, galichon2021unreasonable]. It has gained traction in the last decade as a way of studying complex matching mechanisms, finding applications in the economics of international trade [galichon2021unreasonable, mastrandrea2025optimal], logistics and urban network studies [leite2022urban], demand estimation and pricing [qu2025sinkhorn], decision theory, as well as labor and migration economics [carlier2023sista, hazard2025whom]. We introduce a novel notion of fairness for optimal transport that allows a central planner—such as a government or regulatory body—to specify the probability that individuals from one subgroup are matched with individuals from another subgroup. This flexible definition can enforce a variety of constraints, including limited homogamy (reducing the likelihood that individuals with similar sensitive attributes are matched), matching quotas (ensuring that $p$ percent of a minority subgroup is matched to another group), or broader goals of social diversity. Specifically, our contributions are as follows.
  1. 1.
    We first provide a theoretical framework for fair optimal transport, introducing a formal and novel definition. Building on this framework, we present the FairSinkhorn algorithm for computing optimal transport plans under exact fairness constraints.
  2. 2.
    We design two relaxed approaches. The first one penalizes the optimal transport objective with a convex penalty that promotes fair transport plans. This formulation yields a convex optimization problem in which the utility–fairness trade-off can be easily controlled by adjusting the fairness penalty parameter. The second approach consists in learning a cost function such that the resulting transport plan is fair. It has the advantage of enabling to easily match new samples.
  3. 3.
    For the first approach, we provide a novel sample complexity bound for penalized entropic optimal transport, which we believe to be of independent interest. For the cost learning approach, we provide a bound on the expected fairness deviation when using finite sample to compute the transport plan with the learned cost function.
  4. 4.
    Lastly, we investigate the empirical performance of the proposed approaches through a series of numerical experiments.
Our code will be made public upon publication. Throughout this article, all proofs are deferred to the appendix.
2 Related Work
Fairness in Matching Mechanisms.
Since the pioneering work of gale1962college on the stable marriage problem, a series of works have provided theoretical studies of the fairness in matching mechanisms. A first line of research focuses on the stability of GS-like algorithms under constraints akin to individual fairness, ensuring that similar individuals receive similar outcomes by the decision-making process. For instance, karni2021fairness analyze the Gale-Shapley (GS) algorithm through the lens of preference informed individual fairness. This criterion allows for small deviations from individual fairness, which requires that similar individuals should be treated similarly. They devise a fair variant of the GS algorithm that ensure both approximately fair and stable matchings. Similarly, devic2023fairness analyze merit-based fairness of one-to-one matching mechanisms, a notion close to individual fairness, under partial uncertainty over individual’s merit. A second line of research introduces metrics of fairness closer to group fairness, with a specific focus on kidney paired donation programs, a singular matching problem subject to domain specific constraints — in these programs, donors and patients must be matched by forming cycles due to legal constraints, which leads to a reformulation of the matching problem as a constrained integer programming problem. In their seminal work, ashlagi2014free unveil an unfair market collapse phenomenon in kidney pair donation programs: instead of introducing all patients to the matching system, hospitals internally operate easy matches i.e. do not participate fully in the exchange mechanism. This leads to a loss in transplantation opportunities for highly-sensitized patients that are hard to match. dickerson2014price and following work by zhang2025fairness also consider the case of kidney paired donations programs. Their definition of fairness enforces a minimal amount of program participation for the group of highly-sensitized patients. More recently, lobo2025fair consider a combinatorial multimatching problem under uncertainty: in their setup, the goal is to match agents and resources without exact knowledge of their pairwise valuations. Their definition of fairness relies on the average utility within predefined groups, which is closest to our definition among previous work.
Our contribution differentiates from these works in two aspects. First, we do not consider iterative matching mechanisms such as the GS algorithm: our work focuses on matching through (entropic) optimal transport. Secondly, our definition of fairness is different from previously introduced definitions, and highly modular to accommodate fairness preferences of a central planner. Instead of requiring individual fairness, building on individual rankings over possible assignements or participation constraints, we consider a matching problem between two populations (f.i. students and universities) further structured by a sensitive attribute (f.i. privileged and unprivileged students, and elite and non-elite universities). We enforce matching quotas given by a central planner, that specify for each subgroup the proportion of individuals that should be matched to subgroups in the second population.
Fairness and Optimal Transport.
Recent work has drawn many fruitful connection between fairness and optimal transport, primarily by using OT as a tool to obtain or characterize fair learning algorithms [gordaliza2019obtaining, gouic2020projection, chzhen2020fair, DBLP:conf/aaai/ChiappaJSPJA20, DBLP:conf/nips/BuylB22, hu2023fairness, DBLP:conf/icml/XianY023, chowdhary2024fairness, xiong2024fair, divol2024demographic]. A central insight of this line of work is that learning an optimal fair predictor can be formulated as a Wasserstein barycenter problem [gouic2020projection, chzhen2020fair, divol2024demographic]. This perspective is orthogonal to ours: OT serves as an auxiliary mechanism for analyzing and enforcing fairness in downstream prediction tasks, whereas our goal is to study and enforce fairness of the optimal transport itself.
Closer in spirit to our work is the recent article by nguyen2024towards, which imposes fairness constraints directly on a Wasserstein barycenter by requiring approximate equality of sliced-Wasserstein distances SW [nadjahi2020statistical] to multiple marginals $\mu_{1},\dots,\mu_{K}$:
$\min\limits_{\mu}\sum\limits_{k=1}^{K}\textnormal{{SW}}(\mu,\mu_{k})\textnormal{ s.t }\sum\limits_{i=1}^{K}\sum\limits_{j=1}^{K}\big\lvert\textnormal{{SW}}(\mu,\mu_{i})-\textnormal{{SW}}(\mu,\mu_{j})\big\lvert\leq\varepsilon.$
Their approach, however, does not consider sensitive attributes or demographic groups, focusing instead on balancing distances across marginals. In contrast, we study the problem of computing transport plans that satisfy explicit mass constraints between groups defined by sensitive attributes, a setting that, to the best of our knowledge, has not been previously considered.
Constrained and Statistical Optimal Transport.
Our work borrows from recent developments in constrained optimal transport, which seeks to enforce structural properties on the transport plan [courty2016optimal, blondel2018smooth, paty2019subspace, liu2022sparsity]; for instance, korman2015optimal analyze a variant of optimal transport in which the amount of mass that can be transported between two units is upper bounded. Specifically, we build on rakotomamonjy2015generalized, genevay2019sample and rigollet2022sample to derive finite sample guarantees for a penalized optimal transport problem, and to bound the deviations of fairness with learned cost functions.
3 Formalizing Group Fairness in Optimal Transport
3.1 Optimal Transport
Entropic Optimal Transport.
We consider the entropic regularized optimal transport problem [peyre2019computational, genevay2019sample, feydy2019interpolating, pooladian2021entropic, keriven2022entropic] between two probability distributions $\mu$ and $\eta$ with ground cost $c$:
$\mathbf{W}_{c,\varepsilon}(\mu,\eta):=\min\limits_{\pi\in\Pi(\mu,\eta)}\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\,d\pi(x,y)+\varepsilon\mathbf{KL}\big(\pi|\mu\otimes\eta\big),$
where the minimum is taken over couplings $\Pi(\mu,\eta)$ of $\mu$ and $\eta$ and where $\mathbf{KL}$ denotes the Kullback-Leibler (KL) divergence [CSZISAR_SUBDIFF]. We focus on the entropic formulation because the regularization ensures uniqueness and smoothness of the optimal coupling, which is crucial for efficient computation and gradient-based optimization [peyre2019computational]. With a slight abuse of notation, we will often omit the explicit dependence on $\mu$ and $\eta$ when clear from context, writing, for example, $\mathbf{W}_{c,\varepsilon}$ instead of $\mathbf{W}_{c,\varepsilon}(\mu,\eta)$.
In the finite-sample setting, when $\mu$ and $\eta$ are sums of Dirac measures i.e. if $\mu=1/n\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ and $\eta=1/m\sum_{j=1}^{m}\delta_{\mathbf{y}_{j}}$, the problem above reduces to
$\mathbf{W}_{\mathbf{C},\varepsilon}:=\min\limits_{\mathbf{\Pi}\in\Pi}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{m}\mathbf{\Pi}_{ij}\mathbf{C}_{ij}+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)\kern 5.0pt,$ (1)
where $\Pi:=\big\{\mathbf{\Pi}\in\mathbb{R}_{+}^{n\times m}\,|\,\mathbf{\Pi}\mathbf{1}_{m}=1/n\mathbf{1}_{n},\,\mathbf{\Pi}^{\top}\mathbf{1}_{n}=1/m\mathbf{1}_{m}\big\}$, $\mathbf{KL}\big(\mathbf{\Pi}\big):=\sum_{ij}\mathbf{\Pi}_{ij}\log\mathbf{\Pi}_{ij}$ and $\mathbf{C}\in\mathbb{R}^{n\times m}$ with $\mathbf{C}_{ij}:=c(x_{i},y_{j})$. The KL divergence is strictly convex, and $\Pi$ is a nonempty and bounded set. Therefore, Problem (1) admits a unique minimizer, which we denote by $\mathbf{\Pi}_{\varepsilon}(\mathbf{C})$.
3.2 Fair Optimal Transport
We consider two distributions $\mu\in\mathcal{P}(\mathcal{X}\times\mathcal{S})$ and $\eta\in\mathcal{P}(\mathcal{Y}\times\mathcal{W})$ to be matched, where $\mathcal{X}$ and $\mathcal{Y}$ represent feature spaces while $\mathcal{S}$ and $\mathcal{W}$ correspond to sensitive attributes (e.g., gender, ethnicity or age) defining groups of entities in $\mu$ and $\eta$ respectively. The sets $\mathcal{W}$ and $\mathcal{S}$ are assumed to be finite, and are identified, respectively, with $\{1,\dots,K_{w}\}$ and $\{1,\dots,K_{s}\}$. We first make an assumption on the support of the measures, which is needed to leverage several recent results in statistical optimal transport [rigollet2022sample].
{boxasm}
The measures $\mu$ and $\eta$ are compactly supported.
We let $p_{s}=\mathbb{P}(S=s)$ and $q_{w}=\mathbb{P}(W=w)$ the probability of each group $i$ in $\mu$ and $j$ in $\eta$, respectively, and denote by $\mathbf{p}$ and $\mathbf{q}$ the resulting distributions over $\mathcal{S}$ and $\mathcal{W}$.
Fairness targets.
Figure 1: Summary of the fairness issues arising in optimal transport when sensitive features are highly correlated with features used for cost computations. On the left of the figure, we show an illustrative setting in which we wish to match two populations (circles and triangles). Group membership is highly correlated with the individual’s location in the $(x,y)$ plane. This results in a block sparse transport plan, displayed on the right.
Optimal transport imposes structure on the coupling of two probability distributions through the cost function, making it more likely to match points whose transportation cost to one another is low. While this inductive bias is often desirable, it can lead to undesirable outcomes when features in $\mathcal{X}$ and $\mathcal{Y}$ are correlated with the sensitive attributes in $\mathcal{S}$ and $\mathcal{W}$, as the resulting coupling may reflect, preserve, or amplify disparities between the corresponding groups, as illustrated in the following example.
Example 3.1.
Consider a school assignment system, in which students are characterized by their geographic position $X\in\mathbb{R}^{2}$ and a social status $S\in\{\textnormal{high},\textnormal{low}\}$. Similarly, schools are characterized by geographic positions $Y\in\mathbb{R}^{2}$ and a prestige level $W\in\{\textnormal{elite},\textnormal{non-elite}\}$. Suppose that students with $S=\textnormal{high}$ are more likely to live near elite schools, which are often situated in privileged neighborhoods, while students with low social status tend to live near non-elite schools. In this setting, assigning students to schools using optimal transport with Euclidean cost—minimizing the distance between students and schools—will result in highly segregated schools.
In this example, the block-sparse structure of the optimal transport plan can be seen as a source of unfairness, as the matching will be highly correlated with the social status of the students and the elitist nature of the schools.
To address such issues and define a fair transport plan, we are given a fairness target $\mathbf{F}$, which is a $K_{s}\times K_{w}$ matrix that specifies, for each pair of groups $(s,w)\in\mathcal{S}\times\mathcal{W}$, the desired probability of matching members of group $s$ with group $w$. To be valid, the matrix $\mathbf{F}$ should itself be a coupling between $\mathbf{p}$ and $\mathbf{q}$: it should be non-negative and satisfy the constraints
$\sum\limits_{w=1}^{K_{w}}\mathbf{F}_{sw}=p_{s}\textnormal{ and }\sum\limits_{s=1}^{K_{s}}\mathbf{F}_{sw}=q_{w}$
for all $(s,w)$. We hence write $\Pi(\mathbf{p},\mathbf{q})$ the set of fairness targets.
Example 3.2.
Consider the segregated schooling system introduced in Example 3.1. Let $p:=\mathbb{P}(S=\textnormal{low})$ and $q:=\mathbb{P}(W=\textnormal{non-elite})$. A newly appointed city administrator wishes to strongly limit social homogamy within public schools, and hence requires $60\%$ of student with low social status to be matched to elite schools. The fairness target can be written as
$\mathbf{F}=\begin{bmatrix}0.4\times p&0.6\times p\\ q-0.4\times p&1-q-0.6\times p&\end{bmatrix}$.
Then, we define a coupling to be $\mathbf{F}$-fair if the amount of mass moved from group $s$ to group $w$ matches the fairness target $\mathbf{F}_{sw}$.
Definition 3.3 ($\mathbf{F}$-Fair coupling).
Let $\mu\in\mathcal{P}(\mathcal{X}\times\mathcal{S})$ and $\eta\in\mathcal{P}(\mathcal{Y}\times\mathcal{W})$. We define the $\mathbf{F}$-fair transport plan between $\mu$ and $\eta$ as
$\pi^{\star}_{c,\varepsilon,\textnormal{fair}}\in\operatorname*{arg\,min}\limits_{\pi\in\Pi_{\textnormal{fair}}(\mu,\eta)}\int c(x,y)\,d\pi(x,y)+\varepsilon\mathbf{KL}(\pi|\mu\otimes\eta)\kern 5.0pt.$ (2)
where the set of fair couplings $\Pi_{\textnormal{fair}}(\mu,\eta)$ is defined as
$\Pi_{\textnormal{fair}}(\mu,\eta):=\Big\{\pi\in\Pi(\mu,\eta)\mid\forall(s,w)\in\mathcal{S}\times\mathcal{W},\,\pi_{SW}(s,w)=\mathbf{F}_{sw}\Big\}$
where $\pi_{SW}$ is the coupling on $\mathcal{S}\times\mathcal{W}$ that is obtained from $\pi$ by marginalizing over $x$ and $y$, that is
$\pi_{SW}(s,w):$ $=\pi(\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\})=\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi(x,u,y,t)\kern 5.0pt.$
In the finite-sample case, where we have access to two datasets $(\mathbf{x}_{i},\mathbf{s}_{i})_{i=1}^{n}\in\mathcal{X}\times\mathcal{S}$ and $(\mathbf{y}_{i},\mathbf{w}_{i})_{i=1}^{m}\in\mathcal{Y}\times\mathcal{W}$ drawn i.i.d. from $\mu$ and $\eta$ respectively, the fair optimal transport problem writes
$\begin{aligned} &\min\limits_{\mathbf{\Pi}\in\Pi}\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{C}\big]+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)\\ \textnormal{s.t.}\,\,&\forall(s,w)\in\mathcal{S}\times\mathcal{W},\,\,\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{B}_{sw}\big]=\sum\limits_{i|\mathbf{s}_{i}=s}\sum\limits_{j|\mathbf{w}_{i}=w}\mathbf{\Pi}_{ij}=\mathbf{F}_{sw}\kern 5.0pt\end{aligned}$ (3)
where $\mathbf{B}_{sw}:=\big(\mathds{1}_{\mathbf{s}_{i}=s}\mathds{1}_{\mathbf{w}_{j}=w}\big)_{i,j}\in\{0,1\}^{n\times m}$. This definition may also be extended to include individual specific weights. The following proposition ensures that this problem is well defined.
{boxprop}
Under Assumption 3.2, for any $\mathbf{F}\in\Pi(\mathbf{p},\mathbf{q})$, there exists a unique fair optimal transport plan.
3.3 Solving Fair Optimal Transport
[…]
{boxprop}
There exists $\mathbf{f}\in\mathbb{R}^{n}$, $\mathbf{g}\in\mathbb{R}^{m}$ and $\mathbf{h}=[h_{sw}]_{sw}\in\mathbb{R}^{K_{s}\times K_{w}}$ such that the solution to Problem (LABEL:eq:pure_empirical_fairness) has the form
$\mathbf{\Pi}=\textnormal{diag}\big(e^{\mathbf{f}/\varepsilon}\big)\big(\mathbf{K}\odot\mathbf{H}\big)\textnormal{diag}\big(e^{\mathbf{g}/\varepsilon}\big)\kern 5.0pt,$
where
$\mathbf{K}:=\big[e^{-\mathbf{C}_{ij}/\varepsilon}\big]_{ij}\textnormal{ and }\mathbf{H}:=\sum_{sw}e^{h_{sw}/\varepsilon}\mathbf{B}_{sw}\kern 5.0pt.$
This result mirrors the foundational insight behind the Sinkhorn algorithm and shows that the fairness constraint can be incorporated via a simple additional projection step. This leads to our modified Sinkhorn algorithm, which we call FairSinkhorn, presented in Algorithm˜1 in which the function $\Phi$ is defined as
$\Phi(\mathbf{u},\mathbf{v}):=\bigg(\textnormal{Tr}\Big[\Big(\big[\mathbf{u}_{i}\mathbf{v}_{j}\big]_{ij}\odot\mathbf{B}_{sw}\Big)\big(\mathbf{K}\odot\mathbf{B}_{sw}\big)^{\top}\Big]\bigg)_{sw}.$
The supplementary steps with respect to the original Sinkhorn algorithm are highlighted in blue.
1:Inputs: Cost $\mathbf{C}\in\mathbb{R}^{n\times m}$, $\varepsilon>0$, fairness target $\mathbf{F}=(\mathbf{F}_{sw})_{sw}\in[0,1]^{K_{s}\times K_{w}}$, number of iterations $T$, initializations $\mathbf{u}^{(0)}\in\mathbb{R}^{n}$, $\mathbf{v}^{(0)}\in\mathbb{R}^{m}$, and $\mathbf{L}^{(0)}\in\mathbb{R}^{K_{s}\times K_{w}}$.
[…]
4:for $t=0,\dots,T-1$ do
5:  $\mathbf{u}^{(t+1)}\leftarrow\mathbf{a}\oslash\Big((\mathbf{K}\odot\mathbf{T}^{(t)})\mathbf{v}^{(t)}\Big)$
6:  $\mathbf{v}^{(t+1)}\leftarrow\mathbf{b}\oslash\Big((\mathbf{K}\odot\mathbf{T}^{(t)})^{\top}\mathbf{u}^{(t+1)}\Big)$
7:  $\mathbf{L}^{(t+1)}\leftarrow\mathbf{F}\oslash\Phi(\mathbf{u}^{(t+1)},\mathbf{v}^{(t+1)})$
8:  $\mathbf{T}^{(t+1)}\leftarrow\sum_{sw}\ell^{(t+1)}_{sw}\mathbf{B}_{sw}$ where $\ell^{(t+1)}_{sw}=\big(\mathbf{L}^{(t+1)}\big)_{sw}$
9:end for
10:return $\mathbf{\Pi}=\textnormal{diag}\big(\mathbf{u}^{(T)}\big)\big(\mathbf{K}\odot\mathbf{T}^{(T)}\big)\textnormal{diag}\big(\mathbf{v}^{(T)}\big)$
Algorithm 1 FairSinkhorn algorithm to solve (LABEL:eq:pure_empirical_fairness)
4 Two Strategies for Approximately Fair OT
While the previous section showed that exact fairness can be enforced via a projection step in FairSinkhorn, strictly satisfying fairness constraints can substantially increase transport cost, sometimes beyond what is acceptable in practice. This price of fairness has been studied in operations research for resource allocation problems [bertsimas2011price] and also arises in supervised learning, where enforcing exact group fairness constraints may require sacrificing predictive accuracy [zhao2019inherent]. Motivated by these observations, it is often preferable in practice to allow some tolerance, trading off slight deviations from fairness against overall efficiency.
In this section, we implement this idea by replacing the linear fairness constraints in (LABEL:eq:pure_empirical_fairness) with a relaxed condition
$\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}):=\sum_{(s,w)\in\mathcal{S}\times\mathcal{W}}\bigg(\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{B}_{sw}\big]-\mathbf{F}_{sw}\bigg)^{2}\leq\rho,$ (4)
where $\rho\geq 0$ is a tolerance level. This formulation allows the transport plan to deviate slightly from the exact fairness target while still controlling the total fairness violation. To find optimal transport plans under this relaxed constraint, we propose two strategies.
4.1 Fairness-Penalized OT
Our first strategy is a direct penalization approach. We incorporate the $\rho$-relaxed fairness constraint (4) into the optimal transport problem by introducing a Lagrange multiplier, yielding the penalized objective
$\min_{\mathbf{\Pi}\in\Pi}\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{C}\big]+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)+\lambda\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi})\kern 5.0pt,$ (5)
where $\lambda>0$ controls the strength of the fairness penalty. Importantly, since $\mathcal{L}_{\mathbf{F}}$ is convex, the objective remains strongly convex, ensuring a unique minimizer. This penalized entropic optimal transport problem can be solved efficiently using a generalized conditional gradient algorithm [rakotomamonjy2015generalized]. At each iteration, the algorithm linearizes the fairness regularization around the current iterate and solves a subproblem over the feasible set $\Pi$ to find a search direction. A line search then determines a convex combination of the current iterate and this direction, guaranteeing descent. In practice, the subproblem at iteration $t$ corresponds to an entropic OT problem with the modified cost $\mathbf{C}+\nabla\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}^{t})$ which is solved efficiently with the Sinkhorn algorithm. This makes the algorithm computationally practical even for large-scale problems.
Sample Complexity of Fairness-Penalized OT.
An important contribution of our work is to establish a sample complexity bound for the above fairness-penalized entropic optimal transport cost, by building upon results from entropic optimal transport [genevay2019sample, mena2019statistical]. To formalize this, note that the optimization problem in (5) can be defined for arbitrary measures $\mu\in\mathcal{P}(\mathcal{X}\times\mathcal{S})$ and $\eta\in\mathcal{P}(\mathcal{Y}\times\mathcal{W})$ as (see Section˜B.2 for a formal definition)
$m^{\star}(\mu,\eta):=\min_{\pi\in\Pi(\mu,\eta)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}c(x,y)\,d\pi(x,y)+\varepsilon\mathbf{KL}\big(\pi||\mu\otimes\eta\big)+\lambda\mathcal{L}_{\mathbf{F}}(\pi)\kern 5.0pt.$
Let $\mu_{n}=\sum_{i=1}^{n}\delta_{x_{i}}$ and $\eta_{n}=\sum_{i=1}^{n}\delta_{y_{i}}$ where $(x_{i})_{1\leq i\leq n}$ and $(y_{i})_{1\leq i\leq n}$ are i.i.d. samples drawn from $\mu$ and $\eta$. In Section˜4.1, we quantify how well we can estimate $m^{\star}(\mu,\eta)$ with $m^{\star}(\mu_{n},\eta_{n})$ in function of the sample size $n$. Similar to [rigollet2022sample] and [genevay2019sample], our proof relies on the fact that the measures are compactly supported. We also require the following assumption on the ground cost function.
{boxasm}
The cost function $c:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{+}$ is infinitely differentiable.
{boxthm}
Assume for simplicity that $m=n$. Under Assumptions 3.2 and 4.1, we have
$\mathbb{E}\big\lvert m^{\star}(\mu_{n},\eta_{n})-m^{\star}(\mu,\eta)\big\lvert\lesssim\frac{1}{\sqrt{n}}\kern 5.0pt.$
Compared to existing sample complexity bounds for the standard (non-penalized) entropic OT problem [genevay2019sample, mena2019statistical], our result attains the same scaling with the number of samples $n$, demonstrating that introducing the fairness penalty does not worsen the statistical efficiency of the estimator. The proof can be extended to generic convex penalties at the price of a supplementary $\log(n)$ factor.
Proof sketch for Section˜4.1..
$Y_{n}-m^{\star}(\mu,\eta)\leq m^{\star}(\mu_{n},\eta_{n})-m^{\star}(\mu,\eta)\leq Z_{n}-m^{\star}(\mu,\eta)\kern 5.0pt.$
Lower bound. We obtain $Y_{n}$ through linearization of the convex penalty, which is guaranteed to yield a proper lower bound through the connection between the optimality conditions of our penalized problem and its linearized counterpart [rakotomamonjy2015generalized]. This linearized problem may then be cast as an entropic optimal transport problem with modified cost, allowing us to leverage standard results on the sample complexity of entropic optimal transport [genevay2019entropy, mena2019statistical].
Upper bound. To obtain $Z_{n}$ we consider $\hat{\pi}_{n}^{\star}$, the minimizer of the linearized problem, and insert it in the fairness-penalized loss function, that is we consider
$m^{\star}(\mu_{n},\eta_{n})-m^{\star}(\mu,\eta)\leq\langle c,\hat{\pi}_{n}^{\star}\rangle+\varepsilon\textbf{KL}(\hat{\pi}_{n}^{\star}|\mu_{n}\otimes\eta_{n})+\lambda\mathcal{L}_{F}(\hat{\pi}_{n}^{\star})-m^{\star}(\mu,\eta),$ (6)
where the upper bound follows from the fact that $\hat{\pi}_{n}^{\star}$ satisfies the constraints defining $m^{\star}(\mu_{n},\eta_{n})$. To establish that, in expected value, the absolute value of (6) vanishes at the rate of $1/\sqrt{n}$, we again make use of the connection between the optimality conditions of our penalized problem and its linearized counterpart [rakotomamonjy2015generalized] to reduce the problem to the sample complexity of entropic optimal transport. We handle the part $\lambda\mathcal{L}_{F}(\hat{\pi}_{n}^{\star})-m^{\star}(\mu,\eta)$ by considering a generalization of Theorem 6 from [rigollet2022sample]. ∎
4.2 Fair OT via Cost Learning
The key idea behind our second strategy is to learn a ground cost function $c_{\theta}$ so that the entropic optimal transport (OT) solution $\mathbf{\Pi}(c_{\theta})$ induced by this cost minimizes the fairness violation. In other words, rather than modifying the transport plan directly, we reshape the geometry of the OT problem itself. To control deviations from the original OT problem, we require the learned cost to remain close to a user-chosen baseline cost $c_{\textnormal{base}}$ that encodes prior knowledge or domain-specific structure. This perspective naturally leads to a bilevel optimization problem, where the outer level adjusts the cost parameters and the inner level solves an OT problem whose solution depends on those parameters. Formally, we consider the bilevel problem
$\min\limits_{\theta\in\Theta}\mathcal{L}_{\mathbf{F}}\big(\mathbf{\Pi}_{\varepsilon}(c_{\theta})\big)+\frac{1}{\lambda}\mathscr{D}(c_{\theta},c_{\textnormal{base}})\,\,\textnormal{ s.t. }\,\,\mathbf{\Pi}_{\varepsilon}(c_{\theta})=\operatorname*{arg\,min}\limits_{\mathbf{\Pi}\in\Pi}\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{C}_{\theta}\big]+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)\kern 5.0pt,$ (7)
where $(c_{\theta})_{\theta\in\Theta}$ is a parameterized family of cost functions, $\mathbf{C}_{\theta}:=\big[c_{\theta}(\mathbf{x}_{i},\mathbf{y}_{j})\big]_{ij}$, $c_{\textnormal{base}}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}^{+}$ is a baseline cost function, and $\mathscr{D}$ is a discrepancy measure between cost functions that encourages the learned cost to remain close to the baseline.
Remark 4.1.
This approach is related to prior work on cost learning in optimal transport [JMLR:v20:18-700], which aims to infer a cost function that explains observed matchings or couplings. Here, however, the objective is different: instead of fitting a cost to reproduce a given transport plan, we learn a cost whose induced OT solution optimizes a downstream criterion, namely fairness.
Problem (7) is well posed because the inner problem is an entropic OT problem, which is strongly convex over the set of couplings $\Pi$ and therefore has a unique solution $\mathbf{\Pi}_{\varepsilon}(c_{\theta})$ for any choice of $\theta$. This uniqueness ensures that the bilevel objective is a well-defined function of the cost parameters. In practice, the bilevel problem can be optimized using gradient-based methods: one can differentiate through the iterative solver (iterative differentiation) [Bolte2023one_step_diff, Franceschi2017ForwardReverse, Maclaurin2015, Pauwels2023], or use (approximate) implicit differentiation to approximate the gradient of the bilevel objective [Dagreou2022SABA, Dagreou2024SRBA, Eisenberger2022, Ghadimi2018].
Our approach can accommodate a wide range of parameterizations for $c_{\theta}$. Two concrete examples are provided below, with the Euclidean distance serving as a natural baseline in both cases.
Example 4.1 (Mahalanobis cost).
For a PSD matrix $\mathbf{M}$, the Mahalanobis cost $c_{\mathbf{M}}$ is defined by $c_{\mathbf{M}}(x,y)=(x-y)^{\top}\mathbf{M}(x-y)$. Intuitively, this corresponds to applying a linear transformation to $x$ and $y$ before computing the squared Euclidean distance. A key advantage of this parametrization is its interpretability: the learned matrix $\mathbf{M}$ directly reveals which directions or features in the data are emphasized or downweighted in the transport cost.
Example 4.2 (Neural cost).
A more flexible parametrization uses a neural network $\phi_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ to embed the inputs, defining the cost as $c_{\theta}(x,y)=\lVert\phi_{\theta}(x)-\phi_{\theta}(y)\rVert^{2}_{2}$. The network can be trained from scratch or initialized from a pretrained model and fine-tuned for the specific OT task, enabling highly expressive cost functions that capture complex relationships between $x$ and $y$.
Fairness on Unseen Samples.
A key advantage of our cost learning approach is that it produces a reusable cost function, which can be directly applied to match new samples. This naturally raises the question of whether transport plans obtained by solving an optimal transport problem with the learned cost on previously unseen data will maintain a comparable level of fairness.
To address this question, we provide a deviation bound on the fairness of the transport plan that holds uniformly over all cost functions in the parameterized family. Let $\pi_{\varepsilon}^{\star}(c_{\theta})$ denote the optimal transport plan on two measures $\mu$ and $\eta$. Our goal is to control the difference between the fairness of this plan, defined as
$\mathcal{L}_{\mathbf{F}}\big(\pi_{\varepsilon}^{\star}(c_{\theta})\big):=\sum\limits_{s=1}^{K_{s}}\sum\limits_{w=1}^{K_{w}}\bigg(\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi_{\varepsilon}^{\star}(c_{\theta})(x,u,y,t)-\mathbf{F}_{sv}\bigg)^{2}\kern 5.0pt,$
and the fairness $\mathcal{L}_{\mathbf{F}}\big(\mathbf{\Pi}(c_{\theta})\big)$ of the optimal transport plan $\mathbf{\Pi}(c_{\theta})$ computed on finite-sample versions $\mu_{n}$ and $\eta_{m}$. We will require the following assumption to hold.
{boxasm}
There exists a constant $R_{\Theta}>0$ such that
$\sup\limits_{\theta\in\Theta}\,\,\lVert c_{\theta}\rVert_{\infty}:=\sup\limits_{\theta\in\Theta}\quad\max\limits_{\begin{subarray}{c}x\in\textnormal{supp}(\mu)\\ \,y\in\textnormal{supp}(\eta)\end{subarray}}c_{\theta}(x,y)<R_{\Theta}.$
This assumption is mildly restrictive. For instance, it holds if $\Theta$ is bounded and the function $(\theta,x,y)\mapsto c_{\theta}(x,y)$ is continuous on $\Theta\times\mathcal{X}\times\mathcal{Y}$. For instance, for the Mahalanobis cost function, the parameter space $\Theta$ is a subset of the space of positive definite matrices, and we have
$R_{\Theta}=\max\limits_{\mathbf{M}\in\Theta}\lVert\mathbf{M}\rVert_{\textnormal{F}}\max\limits_{\begin{subarray}{c}x\in\textnormal{supp}(\mu)\\ \,y\in\textnormal{supp}(\eta)\end{subarray}}\lVert x-y\rVert_{2}$
which is finite under Assumption 3.2 as long $\Theta$ is bounded. Similarly, for the neural cost model, provided that the neural network is $C_{\Theta}$-Lipschitz, we have
$R_{\Theta}=C_{\Theta}\max\limits_{\begin{subarray}{c}x\in\textnormal{supp}(\mu)\\ \,y\in\textnormal{supp}(\eta)\end{subarray}}\lVert x-y\rVert_{2}.$
{boxthm}
[Deviation bound for fairness] Assume for simplicity that $n=m$. Under Assumptions 3.2 and 4.2, it holds
$\sup\limits_{\theta\in\Theta}\mathbb{E}\big[\big\lvert\mathcal{L}_{\mathbf{F}}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})\big)-\mathcal{L}_{\mathbf{F}}\big(\pi^{\star}_{\epsilon}(c_{\theta})\big)\big\lvert\big]\lesssim\frac{\exp\big(5R_{\Theta}\varepsilon^{-1}\big)}{\sqrt{n}}.$
Proof sketch for Section˜4.2.
We generalize rigollet2022sample for any bounded cost $c_{\theta}$ by proving that for any test function $\varphi\in L^{\infty}(\mu\otimes\eta)$ and for any $t>0$, we have with probability $1-18e^{-t^{2}}$
$\big\lvert\int\varphi d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert\lesssim\frac{\exp\big(5R_{\Theta}\epsilon^{-1}\big)\lVert\varphi\rVert_{L^{\infty}(\mu\otimes\eta)}\cdot t}{\sqrt{n}}$
which yield the following bound in expectation
$\mathbb{E}\big[\big\lvert\int\varphi d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert\big]\lesssim\frac{\exp\big(5R_{\Theta}\epsilon^{-1}\big)\lVert\varphi\rVert_{L^{\infty}(\mu\otimes\eta)}}{\sqrt{n}}.$
Then, using the identity $\lvert a^{2}-b^{2}\rvert\leq\lvert a+b\rvert\lvert a-b\rvert$, we show that
$\lvert\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}_{\epsilon}(c_{\theta}))-\mathcal{L}_{\mathbf{F}}(\pi^{\star}_{\epsilon}(c_{\theta}))\rvert\lesssim\sum_{s=1}^{K_{s}}\sum_{w=1}^{K_{w}}\lvert\textnormal{Tr}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})^{\top}\mathbf{B}_{uv}\big)-\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi^{\star}_{\epsilon}(c_{\theta})(x,u,y,t)\big\lvert$.
For suitable choice of $\varphi_{sw}\in L^{\infty}(\mu\otimes\eta)$, we have for any $s,w\in[K_{s}]\times[K_{w}]$
$\lvert\textnormal{Tr}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})^{\top}\mathbf{B}_{uv}\big)-\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi^{\star}_{\epsilon}(c_{\theta})(x,u,y,t)\big\lvert=\big\lvert\int\varphi_{sw}d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert$
Putting all together yields the result. ∎
This bound provides a uniform control over the expected difference between the fairness loss evaluated on the population and its finite-sample approximation, for any cost function $c_{\theta}$ in the parameterized family. The exponential dependence on $\varepsilon^{-1}$ is unavoidable in the worst case; see rigollet2022sample and altschuler2022asymptotics for related discussions.
Importantly, Section˜4.2 implies that, for sufficiently large training and test samples, the transport plan induced by the learned cost achieves a level of fairness on the test samples similar to that observed on the training samples. In other words, the learned cost can be reused to match new samples with fairness guarantees.
4.3 Comparison of Both Approaches
Convexity.
The penalized OT problem is convex, whereas the cost learning approach involves a bilevel optimization, which is non-convex and may get stuck in local minima.
Flexibility and Interpretability.
Cost learning offers greater flexibility: the baseline cost function and the discrepancy $\mathscr{D}$ can be chosen by the user. Moreover, selecting an appropriate parameterized family can make the learned cost interpretable.
Reusability.
Penalized OT requires solving a new problem for each batch of samples. In contrast, a learned cost can be reused to match new samples via a standard OT problem.
5 Experiments
5.1 Approaches and Metrics
Approaches.
We compare our two relaxed approaches: penalized OT and cost learning. For cost learning, we evaluate two parameterizations: a Mahalanobis cost (˜4.1) and a neural cost (˜4.2), where $\phi_{\theta}$ is a multilayer perceptron (MLP) or a ResNet trained from scratch.
5.2 Approaches and Metrics
Approaches.
We compare our two relaxed approaches: penalized OT and cost learning. For cost learning, we evaluate two parameterizations: a Mahalanobis cost (˜4.1) and a neural cost (˜4.2), where $\phi_{\theta}$ is a multilayer perceptron (MLP) or a ResNet embedding.
Metrics.
Following standard practice in fairness research, we analyze the utility–fairness trade-off by visualizing performance in the two-dimensional (utility, fairness) plane. The Pareto front corresponds to all points that, for any given level of fairness, achieve the lowest level of bias. The Pareto front consists of points that, for a given level of fairness, achieve the lowest possible bias; points on this front are considered optimal, as improving one metric requires trading off the other. Bias is measured relative to the transport plan obtained by standard (non-fair) OT with the same entropic regularization, using two quantities: (a) the KL divergence between transport plans, and (b) the difference in total transport cost. In all experiments, the target $\mathbf{F}$ has all coefficients equal to zero, except for the anti-diagonal entries, which are set to $\frac{1}{KsKw}$.
Implementation.
Our experiments are conducted using the Python Optimal Transport library [flamary2021pot, flamary2024pot], which provides implementations of the Sinkhorn and generalized conditional gradient algorithms. In the cost-learning approach, we use iterative differentiation by backpropagating through the Sinkhorn iterations [Pauwels2023] to compute gradients of the bilevel objective using the PyTorch framework [Paszke2019].
5.3 Comparing Exact and Penalized Approaches
To illustrate the need to move from strict group fairness constraints to relaxed ones, we compare the transport cost and the level of fairness achieved for the transport plans obtained using our FairSinkhorn algorithm, introduced in Section˜3.2, the transport plan obtained using the original Sinkhorn algorithm, and the fairness-penalized optimal transport presented in Section˜4.1, for different values of the fairness penalty. In Fig.˜2, one observes that FairSinkhorn achieves perfect fairness, but with high transport cost. In contrast, Sinkhorn achieves low transport cost with high fairness loss. One observes that adopting the relaxed formulation allows to attain low fairness violations while achieving a better transport cost. This demonstrate that the penalization enables to tune the fairness-utility trade-off of interest.
Figure 2: Comparison of FairSinkhorn , penalized optimal transport, and Sinkhorn on the Gaussian data introduced in Section˜5.4. FairSinkhorn achieves perfect fairness with a high transport cost while Sinkhorn achieves low transport cost with low fairness. The penalized OT interpolates between Sinkhorn and FairSinkhorn .
5.4 Comparing Penalized OT and Cost Learning
Figure 3: Synthetic Gaussian data experiment. Left: 2D gaussians with binary sensitive attribute $S\in\{0,1\}$ and $W\in\{0,1\}$. Right: Utility/fairness tradeoffs. On the top, the utility metric is the KL divergence between the OT plan with and without the fairness constraint in function of the fairness regularization parameter. On the bottom, the difference between the transport cost with fair and unfair transport plan. The penalized OT correspond to our first relaxed approach while de Mahalanobis, MLP and ResNet ones correspond to the cost learning approach.
Figure 4: Transport plan we get using different level of fairness parameter for each approach with the Gaussian data.
We now evaluate both approaches on simulated datasets. We consider two types of data. First, we study mixtures of Gaussians, where each group is assigned to a distinct mixture component (see Fig.˜3, left). Second, we consider a dataset composed of concentric circles (see Fig.˜5, left), which allows us to examine the effect of data geometry. In both settings, the penalized OT approach achieves a better utility–fairness trade-off, benefiting from its convex formulation. Moreover, for the cost-learning approach, we observe that using more expressive cost models enables lower levels of fairness loss to be attained.
Figure 5: Synthetic circular data experiment. Left: 2D gaussians with sensitive attribute $S\in\{0,1,2\}$ and $W\in\{0,1,2\}$. Right: Utility/fairness tradeoffs. On the top, the utility metric is the KL divergence between the OT plan with and without the fairness constraint in function of the fairness regularization parameter. On the bottom, the difference between the transport cost with fair and unfair transport plan. The penalized OT correspond to our first relaxed approach while de Mahalanobis, MLP and ResNet ones correspond to the cost learning approach.
Figure 6: Transport plan we get using different level of fairness parameter for each approach with the circular data.
5.5 Assessing the Reusability of Learned Costs
One of the peaks of our cost learning driven approach is that once learned, the cost function can be called on new samples without any retraining. In contrast, the penalized optimal transport plan cannot be reused, and we need to solve a new transport problem for every new batch of data points.
We quantify the reusability of our bilevel algorithm in the following way. We first learn a Mahalanobis cost function $\widehat{\mathbf{M}}$ on a set of $n_{0}=200$ datapoints with $k_{s}=k_{w}=2$ distinct classes. We subsequently sample $k=50$ new datapoints from the same distribution, and $a)$ solve the penalized OT problem on this fresh dataset $b)$ solve the optimal transport problem using $\widehat{\mathbf{M}}$ to define the cost and compute the fairness of the resulting transport plan. Results are displayed in Figure 7. One observes that using the learned cost allows new data points to be matched with significantly lower inference time than the penalized approach, since the latter requires solving the entire penalized optimal transport problem each time new samples are added to the dataset.
Figure 7: We assess the recyclability of the learned cost function by $a)$ learning a cost function on $n_{0}=200$ samples and $b)$ computing the entropic transport plan on batches of $n_{k}$ new unseen samples. Concurrently to this setup, we train the penalized transport plan from scratch for every new batch. Here, $\varepsilon=2$, $n_{k}=100$, $d=2$ and $k_{s}=k_{w}=2$. We set the penalty for all three approaches to yield comparable levels of fairness, but emphasize that our goal in this experiment is to compare the running times of the different approaches.
6 Conclusion
In this paper, we introduced a novel notion of group fairness for optimal transport. Building on this new notion, we proposed an algorithm that solves entropic regularized optimal transport with perfect fairness, though this can lead to solutions with high transport cost. To achieve a better utility–fairness trade-off, we then propose two relaxed approaches: (i) a convex penalization of the entropic optimal transport objective, and (ii) a cost-learning-based method. For the penalized approach, we derive a sample complexity bound. For the cost learning approach, we bound the expected deviation of the fairness loss when the transport plan is estimated from a finite number of samples drawn from the underlying distributions. Deriving a generalization bound for the cost learning approach is left for future work.
Acknowledgments.
This work was mostly conducted while LB was a member of Inria Montpellier. LB’s work at EPFL is supported by an EPFL AI Center Postdoctoral Fellowship. Parts of the experiments for this article were run on the RCP cluster at EPFL. F.A.’s work was partially supported by a French government grant managed by the Agence Nationale de la Recherche under the France 2030 program, reference SMATCH ANR-22-PESN-0003. We thank the anonymous reviewer and all participants of the Eurips workshop Unifying Perspectives on Learning Biases whose comments improved this manuscript.
Appendix A A Modified Sinkhorn Algorithm for Fair Optimal Transport
The following is an immediate consequence of first order conditions on the Lagrangian of Problem LABEL:eq:pure_empirical_fairness.
{boxprop}
There exists $\mathbf{f}\in\mathbb{R}^{n}$, $\mathbf{g}\in\mathbb{R}^{m}$ and $\mathbf{h}=[h_{sw}]_{sw}\in\mathbb{R}^{K_{s}\times K_{w}}$ such that the solution to Problem LABEL:eq:pure_empirical_fairness has the form
$\mathbf{\Pi}=\textnormal{diag}\big(e^{\mathbf{f}/\varepsilon}\big)\big(\mathbf{K}\odot\mathbf{H}\big)\textnormal{diag}\big(e^{\mathbf{g}/\varepsilon}\big)$
where
$\mathbf{K}:=\big[e^{-\mathbf{C}_{ij}/\varepsilon}\big]_{ij}\textnormal{ and }\mathbf{H}:=\sum_{sw}e^{h_{sw}/\varepsilon}\mathbf{B}_{sw}.$
Proof.
Introducing dual variables $\mathbf{f}\in\mathbb{R}^{n},\mathbf{g}\in\mathbb{R}^{m}$ and $\mathbf{H}=(h_{sw})_{sw}\in\mathbb{R}^{K_{s}\times K_{w}}$, the Lagrangian writes
$\mathcal{E}\big(\mathbf{\Pi},\mathbf{f},\mathbf{g},\mathbf{h}\big)=$ $\textnormal{Tr}[\mathbf{\Pi}^{\top}\mathbf{C}]+\varepsilon\mathbf{KL}(\mathbf{\Pi})-\mathbf{f}^{\top}(\mathbf{\Pi}\mathds{1}_{m}-\mathbf{a})-\mathbf{g}^{\top}(\mathbf{\Pi}^{\top}\mathds{1}_{n}-\mathbf{b})$
- $\sum\limits_{s,w}h_{sw}\Big[\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{B}_{sw}\big]-\mathbf{F}_{sw}\Big].$
First order conditions yield for every $i,j$
$\frac{\partial\mathcal{E}\big(\mathbf{\Pi},\mathbf{f},\mathbf{g},\mathbf{h}\big)}{\partial\mathbf{\Pi}_{ij}}=\mathbf{C}_{ij}+\varepsilon\log\big(\mathbf{\Pi}_{ij}\big)+\varepsilon-\mathbf{f}_{i}-\mathbf{g}_{j}-\sum\limits_{sw}h_{sw}\big[\mathbf{B}_{sw}\big]_{ij}=0$
which we may rewrite as
$\mathbf{\Pi}_{ij}=$ $\exp\big(\mathbf{f}_{i}/\varepsilon\big)\exp\Big(-\varepsilon^{-1}\mathbf{C}_{ij}+\varepsilon^{-1}\sum\limits_{sw}h_{sw}\big[\mathbf{B}_{sw}\big]_{ij}-1\Big)\exp\big(\mathbf{g}_{j}/\varepsilon\big)$
$=$ $\exp\big(\mathbf{f}_{i}/\varepsilon\big)\exp(-\mathbf{C}_{ij}/\varepsilon-1)\prod_{s^{\prime}=1}^{K_{s}}\prod_{w^{\prime}=1}^{K_{w}}\exp\big(h_{s^{\prime}w^{\prime}}[\mathbf{B}_{s^{\prime}w^{\prime}}]_{ij}/\varepsilon\big)\exp\big(\mathbf{g}_{j}/\varepsilon\big)\kern 5.0pt.$
Now, remark that for any $(i,j)\in\{1,\dots,n\}^{2}$, there is only one $(s,w)\in\{1,\dots,K_{s}\}\times\{1,\dots,K_{w}\}$ such that $\mathbf{B}_{sw}]_{i,j}\neq 0$. Therefore, in the product $\prod_{s^{\prime}=1}^{K_{s}}\prod_{w^{\prime}=1}^{K_{w}}\exp\big(h_{s^{\prime}w^{\prime}}[\mathbf{B}_{s^{\prime}w^{\prime}}]_{ij}\big)$, there is only one term distinct from $1$. As a consequence
$\prod_{s^{\prime}=1}^{K_{s}}\prod_{w^{\prime}=1}^{K_{w}}\exp\big(h_{s^{\prime}w^{\prime}}[\mathbf{B}_{s^{\prime}w^{\prime}}]_{ij}/\varepsilon\big)$ $=\prod_{s^{\prime}=1}^{K_{s}}\prod_{w^{\prime}=1}^{K_{w}}\exp\big(h_{s^{\prime}w^{\prime}}\mathds{1}_{\mathbf{s}_{i}=s^{\prime}}\mathds{1}_{\mathbf{w}_{j}=w^{\prime}}/\varepsilon\big)$
$=\sum\limits_{s^{\prime}=1}^{K_{s}}\sum\limits_{w^{\prime}=1}^{K_{w}}\exp\big(h_{s^{\prime}w^{\prime}}/\varepsilon\big)\mathds{1}_{\mathbf{s}_{i}=s^{\prime}}\mathds{1}_{\mathbf{w}_{j}=w^{\prime}}$
$=\sum\limits_{s^{\prime}=1}^{K_{s}}\sum\limits_{w^{\prime}=1}^{K_{w}}\exp\big(h_{s^{\prime}w^{\prime}}/\varepsilon\big)\big[\mathbf{B}_{s^{\prime}w^{\prime}}\big]_{ij}\kern 5.0pt.$
This leads to the matrix form
$\mathbf{\Pi}=\textnormal{diag}\big(e^{\mathbf{f}/\varepsilon}\big)\big(\mathbf{K}\odot\mathbf{H}\big)\textnormal{diag}\big(e^{\mathbf{g}/\varepsilon}\big)$
where
$\mathbf{K}:=e^{-\mathbf{C}/\varepsilon-1}$
$\mathbf{H}:=\sum\limits_{s=1}^{K_{s}}\sum\limits_{w=1}^{K_{w}}\exp({h_{sw}/\varepsilon})\odot\mathbf{B}_{sw}.$
A fair optimal transport plan can hence be obtained using a modified Sinkhorn algorithm which we call FairSinkhorn. Define the map
$\Phi(\mathbf{u},\mathbf{v}):=\bigg(\textnormal{Tr}\Big[\Big(\big[\mathbf{u}_{i}\mathbf{v}_{j}\big]_{ij}\odot\mathbf{B}_{sw}\Big)\big(\mathbf{K}\odot\mathbf{B}_{sw}\big)^{\top}\Big]\bigg)_{sw}.$
The algorithm FairSinkhorn combines projections on the set of marginal constraints with a projection on the set of fairness constraints, highlighted in blue.
Remark 1.
Since FairSinkhorn alternates between projections on sets defined by linear constraints, it is guaranteed to converge [benamou2015iterative].
Appendix B Proofs
B.1 Proof of Section˜3.2
Lemma B.1.
Let $\mu\in\mathcal{P}(\mathcal{X}\times\mathcal{S})$ and $\mathcal{\eta}\in\mathcal{P}(\mathcal{Y}\times\mathcal{W})$. Let $p\in\mathcal{P}(\mathcal{S})$ and $q\in\mathcal{P}(\mathcal{W})$ be obtained from $\mu$ and $\eta$ by marginalizing, respectively, $x$ and $y$, that is,
$p(S=s)$ $=\mu(\mathcal{X}\times\{s\})$ (8)
$q(W=w)$ $=\eta(\mathcal{Y}\times\{w\}).$ (9)
Finally let $F\in\Pi(p,q)$.
There exists $\pi\in\Pi(\mu,\eta)$ such that $\pi$ is $F$-fair, that is
$\pi\big(\mathcal{\mathcal{X}}\times\{s\}\times\mathcal{\mathcal{Y}}\times\{w\}\big)=F(S=s,W=w).$
Proof.
Given measurable sets $A\subseteq\mathcal{X}$ and $B\subseteq\mathcal{Y}$ let $\pi$ be defined as
$\pi\big(A\times\{s\}\times B\times\{w\}\big):=\frac{F(S=s,W=w)\mu(A\times\{s\})\eta(B\times\{w\})}{p(S=s)q(W=w)}$
and let’s check that i) $\pi\in\Pi(\mu,\eta)$ and ii) $\pi$ is $F$-fair:
  • i)
    We want to check that $\pi(A\times\{s\}\times\mathcal{Y}\times\mathcal{W})=\mu(A\times\{s\})$ and similarly for the other marginal. To this end note that
    $\pi(A\times\{s\}\times\mathcal{Y}\times\mathcal{W})$ $=\sum_{w\in\mathcal{W}}\frac{F(S=s,W=w)\mu(A\times\{s\})\eta(\mathcal{Y}\times\{w\})}{p(S=s)q(W=w)}$
    $=\mu(A\times\{s\})\sum_{w\in\mathcal{W}}\frac{F(S=s,W=w)}{p(S=s)},$
    where the second equality follows from (9). To finish note that from $F\in\Pi(p,q)$ follows that $\sum_{w\in\mathcal{W}}F(S=s,W=w)=p(S=s)$. A similar argument shows that $\pi(\mathcal{X}\times\mathcal{S}\times B\times\{w\})=\eta(B\times\{w\})$.
  • ii)
    To see that $\pi$ is $F$-fair note that it is immediate from (8) and (9) that
    $\pi\big(\mathcal{\mathcal{X}}\times\{s\}\times\mathcal{\mathcal{Y}}\times\{w\}\big)$ $=\frac{F(S=s,W=w)\mu(\mathcal{X}\times\{s\})\eta(\mathcal{Y}\times\{w\})}{p(S=s)q(W=w)}$
    $=F(S=s,W=w).$
Proposition B.1.
Assume that $\mu$ and $\eta$ have bounded support. Let $F$ be a coupling of the marginals of $\mu$ and $\eta$ denoted by $p$ and $q$ defined via (8)-(9). Then there exists a unique $F$-fair transport plan.
Proof.
Assume $\mathcal{X}$ and $\mathcal{Y}$ to be compact. Similar to the proof of Theorem 1.4. in [santambrogio2015optimal] we can prove that
$\Lambda:=\big\{\pi\in\mathcal{P}(\mathcal{X}\times\{0,1\}\times\mathcal{Y}\times\{0,1\}):\pi_{SW}=F\big\}$
is compact with respect to the weak topology: Let $\pi_{n}$ be a sequence in $\Lambda$. They are probability measures, so that their mass is 1, and hence they are bounded in the dual of $C(\mathcal{X}\times\{0,1\}\times\mathcal{Y}\times\{0,1\})$. Hence usual weak-$\star$ compactness in dual spaces guarantees the existence of a subsequence $\pi_{n}\rightharpoonup\pi$ converging to a probability $\pi$. We just need to check $\pi\in\Lambda$. This may be done by fixing $\phi\in C(\{0,1\}\times\{0,1\})$ and from $\pi_{n}\in\Lambda$ it follows that
$\int\phi\,d\pi_{n}=\int\phi\,d\big[(\pi_{n})_{SW}\big]=\int\phi\,dF=\sum_{s,w}\phi(s,w)F(S=s,W=w).$
Now pass to the limit to obtain
$\int\phi\,d\pi_{SW}=\int\phi\,d\pi=\sum_{s,w}\phi(s,w)F(S=s,W=W)$
which shows that $\pi\in\Lambda$.
To finish, just note that $\Pi_{\text{fair}}^{\mathbf{F}}=\Pi(\mu,\eta)\cap\Lambda$ is the intersection of two compact sets (with respect to weak topology) and Lemma B.1 establishes that it is non-empty. Continuity of the map defining the transport problem is enough to conclude the existence of a minimizer. Uniqueness is then a consequence of the strict convexity of KL. ∎
B.2 Proof of Section˜4.1
Notation.
Let $\mathbf{F}\in\mathbb{R}^{K_{s}\times K_{w}}$ a given a fairness target. We adopt the following notational conventions, $\mathbf{u}\in[0,1]^{K_{s}}$, $\mathbf{u}^{\top}\mathbf{1}_{K_{s}}=1$, and $\mathbf{t}\in[0,1]^{K_{w}}$, $\mathbf{t}^{\top}\mathbf{1}_{K_{w}}=1$, we define the penalization to be
$\xi_{sw}(\mathbf{u},\mathbf{t}):=\mathbf{u}_{s}\mathbf{t}_{w}-\mathbf{F}_{sw},$ (10)
so that given a measure $\pi\in\mathcal{M}(\mathcal{X}\times[K_{s}]\times\mathcal{Y}\times[K_{w}])$, define
$\mathcal{L}_{\mathbf{F}}(\pi)\vcentcolon=\sum_{s,w\in[K_{s}]\times[K_{w}]}\langle\xi_{sw},\pi\rangle^{2}=\sum_{s,w\in[K_{s}]\times[K_{w}]}{\left(\int\xi_{sw}(z,w)\,d\pi(x,z,y,w)\right)}^{2},$
where we use the duality pairing notation and integral notation interchangeably. For any distribution $\mu$, $\eta$ over a set $\mathcal{X}$ such that $\mu$ is absolutely continuous with respect to $\eta$ (i.e., $\mu\ll\eta$), we denote by $\frac{d\mu}{d\eta}$ the Radon-Nikodym derivative of $\mu$ with respect to $\eta$, and
$\mathbf{KL}\big(\mu||\eta\big)\vcentcolon=\int_{\mathcal{X}}\log\left(\frac{d\mu}{d\eta}(x)\right)d\mu(x).$
Given any two measures $\mu\in\mathcal{M}(\mathcal{X}\times[K_{s}])$ and $\eta\in\mathcal{M}(\mathcal{Y}\times[K_{w}])$ define
$m^{\star}(\mu,\eta):=\min_{\pi\in\Pi(\mu,\eta)}\langle c,\pi\rangle+\varepsilon\mathbf{KL}\big(\pi||\mu\otimes\eta\big)+\lambda\mathcal{L}_{\mathbf{F}}(\pi).$ (11)
The goal is to prove the following result
Theorem 4.1.
Let $\mathcal{X}$ and $\mathcal{Y}$ be compact subsets of $\mathbb{R}^{d}$ and let $\mu$ and $\eta$ be probability measures on $\mathcal{X}\times[K_{s}]$ and $\mathcal{Y}\times[K_{w}]$, respectively. Let $(x_{i},s_{i})_{i=1}^{n}$ be $n$ independent and identically distributed (i.i.d.) samples from $\mu$ and let $(y_{j},w_{j})_{j=1}^{n}$ be $n$ i.i.d. samples from $\eta$; assume further that the samples from $\mu$ are independent from those from $\eta$. Finally, suppose that the cost $c$ is $L$-Lipschitz continuous and infinitely differentiable. Then
$\mathbb{E}_{\mu\otimes\eta}|m^{\star}(\mu_{n},\eta_{n})-m^{\star}(\mu,\eta)|\leq\mathcal{O}(1/\sqrt{n}^{-1}),$ (12)
Proof.
Let $m^{\star}_{\infty}:=m^{\star}(\mu,\eta)$ and $m^{\star}_{n}:=m^{\star}(\mu_{n},\eta_{n})$, and $\pi_{\infty}^{\star}$ be the minimizer attaining the $m^{\star}_{\infty}$ that is a measure on $\mathcal{X}\times[K_{s}]\times\mathcal{Y}\times[K_{w}]$ with marginals $\mu$ and $\eta$ such that
$m_{\infty}^{\star}=\langle c,\pi^{\star}_{\infty}\rangle+\varepsilon\mathbf{KL}\big(\pi^{\star}_{\infty}||\mu\otimes\eta\big)+\lambda\mathcal{L}_{\mathbf{F}}(\pi^{\star}_{\infty}).$
Let $p^{\star}_{\infty}$ denote the Radon–Nikodym density of $\pi^{\star}_{\infty}$ with respect to $\mu\otimes\eta$ (note that the Kullback–Leibler term forces $\pi^{\star}_{\infty}$ to be absolutely continuous with respect to $\mu\otimes\eta$).
The idea is to “sandwich” the random variable $m_{n}^{\star}-m^{\star}_{\infty}$ between two random variables with expectation upper bounded by $\mathcal{O}(1/\sqrt{n})$, that is, to find random variables $Y_{n}$, $Z_{n}$ such that
$Y_{n}-m_{\infty}^{\star}\leq m_{n}^{\star}-m_{\infty}^{\star}\leq Z_{n}-m_{\infty}^{\star}$ (13)
and
$\mathbb{E}[|Y_{n}-m_{\infty}^{\star}|]\leq\mathcal{O}(1/\sqrt{n}^{-1})\quad\text{ and }\quad\mathbb{E}[|Z_{n}-m_{\infty}^{\star}|]\leq\mathcal{O}(1/\sqrt{n}^{-1}).$
To see that this is enough to establish (12), observe that (13) implies that
$|m_{n}^{\star}-m_{\infty}^{\star}|\leq\max(|Y_{n}-m_{\infty}^{\star}|,|Z_{n}-m_{\infty}^{\star}|)\leq|Y_{n}-m_{\infty}^{\star}|+|Z_{n}-m_{\infty}^{\star}|.$
Lower bound.
To define $Y_{n}$ start by observing that by linearizing the fairness penalization, we can incorporate this linearization into the transport cost, so that the problem is now a proper entropic optimal transport one. This allows us to leverage existing results on the sample complexity of optimal transport. Moreover, the convexity of $\mathcal{L}_{\mathbf{F}}$ implies that this linearization does yield a lower bound. To this end, note that for any measure $\pi\in\mathcal{M}(\mathcal{X}\times[K_{s}]\times\mathcal{Y}\times[K_{w}])$ and any $s,w\in[K_{s}]\times[K_{w}]$, the inequality $0\leq(\langle\xi_{sw},\pi-\pi^{\star}_{\infty}\rangle)^{2}=(\langle\xi_{sw},\pi\rangle)^{2}+(\langle\xi_{sw},\pi_{\infty}^{\star}\rangle)^{2}-2\langle\xi_{sw},\pi\rangle\langle\xi_{sw},\pi^{\star}_{\infty}\rangle,$ implies that
$\mathcal{L}_{\mathbf{F}}(\pi)\geq$ $-\mathcal{L}_{\mathbf{F}}(\pi_{\infty}^{\star})+2\sum_{(s,w)\in[K_{s}]\times[K_{w}]}\big\langle\langle\xi_{sw},\pi^{\star}_{\infty}\rangle\xi_{sw},\pi\big\rangle$
$=$ $\mathcal{L}_{\mathbf{F}}(\pi^{\star}_{\infty})+\big\langle 2\sum_{s,w\in[K_{s}]\times[K_{w}]}\langle\xi_{sw},\pi^{\star}_{\infty}\rangle\xi_{sw},\pi-\pi^{\star}_{\infty}\big\rangle.$
This leads to a lower bound on $m_{n}^{\star}$ given by
$m_{n}^{\star}\geq-\lambda\mathcal{L}_{\mathbf{F}}(\pi_{\infty}^{\star})+\underbrace{\min_{\begin{subarray}{c}\pi_{1}=\mu_{n}\\ \pi_{2}=\eta_{n}\end{subarray}}\langle\overbrace{c+2\lambda\sum_{s,w\in[K_{s}]\times[K_{w}]}\langle\xi_{sw},\pi_{\infty}^{\star}\rangle\xi_{sw}}^{:=\widehat{c}},\pi\rangle+\varepsilon\mathbf{KL}(\pi||\mu_{n}\otimes\eta_{n})}_{:=\hat{m}_{n}^{\star}}=Y_{n}$ (14)
where $\widehat{m}_{n}^{\star}$ is an entropic optimal transport problem with cost $\widehat{c}$. From the sample complexity of optimal transport (see Theorem 18 in [genevay2019entropy]) it follows that
$\mathbb{E}_{\mu\otimes\eta}\big|\widehat{m}_{n}^{\star}-\widehat{m}_{\infty}^{\star}\big|\leq\mathcal{O}(\sqrt{n}^{-1}),$ (15)
with $\widehat{m}^{\star}_{\infty}$ being the population version of $\widehat{m}_{n}^{\star}$, i.e.,
$\widehat{m}^{\star}_{\infty}:=\min_{\begin{subarray}{c}\pi_{1}=\mu\\ \pi_{2}=\eta\end{subarray}}\langle c+2\lambda\sum_{s,w\in[K_{s}]\times[K_{w}]}\langle\xi_{sw},\pi_{\infty}^{\star}\rangle\xi_{sw},\pi\rangle+\varepsilon\mathbf{KL}(\pi||\mu\otimes\eta).$
To finish the lower bound, note that $\mathbb{E}[|Y_{n}-m^{\star}_{\infty}|]\leq\mathcal{O}(\sqrt{n}^{-1})$ follows from (15) provided we show that $m_{\infty}^{\star}=\widehat{m}_{\infty}^{\star}-\lambda\mathcal{L}_{\mathbf{F}}(\pi_{\infty}^{\star})$; this is the content of Lemma B.2 that can be found in Appendix B.2.1.
Upper bound.
Let $\hat{\pi}_{n}^{\star}$ be a minimizer attaining $\hat{m}_{n}^{\star}$ (see (14) for a definition of $\hat{m}_{n}^{\star}$). From the fact that $\hat{\pi}_{n}^{\star}$ satisfies the marginal constraints defining $m^{\star}_{n}$ follows that
$\min_{\pi\in\Pi(\mu_{n},\eta_{n})}\langle c,\pi\rangle+\varepsilon\mathbf{KL}\big(\pi||\mu_{n}\otimes\eta_{n}\big)+\lambda\mathcal{L}_{\mathbf{F}}(\pi)=m_{n}^{\star}\leq\langle c,\hat{\pi}_{n}^{\star}\rangle+\varepsilon\mathbf{KL}\big(\hat{\pi}_{n}^{\star}||\mu_{n}\otimes\eta_{n}\big)+\lambda\mathcal{L}_{\mathbf{F}}(\hat{\pi}_{n}^{\star}):=Z_{n}.$
Let $\hat{c}$ be as in (14) and note that $Z_{n}$ decomposes as
$Z_{n}=\underbrace{\langle\hat{c},\hat{\pi}_{n}^{\star}\rangle+\varepsilon\mathbf{KL}\big(\hat{\pi}_{n}^{\star}||\mu_{n}\otimes\eta_{n}\big)}_{:=Z_{n_{1}}}+\underbrace{\lambda\mathcal{L}_{\mathbf{F}}(\hat{\pi}_{n}^{\star})+\langle c-\hat{c},\hat{\pi}_{n}^{\star}\rangle}_{:=Z_{n_{2}}}.$
Moreover, by definition of $\hat{\pi}_{n}^{\star}$ and $\hat{c}$,
$Z_{n_{1}}=\widehat{m}_{n}^{\star}\quad\text{ and }\quad Z_{n_{2}}=\lambda\mathcal{L}_{\mathbf{F}}(\hat{\pi}_{n}^{\star})-2\lambda\sum_{(s,w)\in[K_{s}]\times[K_{w}]}\big\langle\langle\xi_{sw},\pi^{\star}_{\infty}\rangle\xi_{sw},\hat{\pi}_{n}^{\star}\big\rangle.$
Lemma B.2 shows that $m^{\star}_{\infty}=\widehat{m}_{\infty}^{\star}-\lambda\mathcal{L}_{F}(\pi^{\star}_{\infty}\rangle$ and, similar to the lower bound part of the proof (see (15)), the sample complexity of entropic optimal transport implies that,
$\mathbb{E}\big[\big|Z_{n_{1}}-\widehat{m}_{\infty}^{\star}\big|\big]=\mathbb{E}\big[\big|\widehat{m}_{n}^{\star}-\widehat{m}_{\infty}^{\star}\big|\big]\leq\mathcal{O}(\sqrt{n}^{-1}).$
Consequently, to establish that $\mathbb{E}[|Z_{n}-m^{\star}_{\infty}|]\leq\mathcal{O}(1/\sqrt{n})$ it is enough to show that
$\mathbb{E}\big[\big|Z_{n_{2}}-\big(-\lambda\mathcal{L}_{F}(\pi_{\infty}^{\star})\big)\big|\big]\leq\mathcal{O}(1/\sqrt{n}).$ (16)
To this end, note that
$\lambda\mathcal{L}_{\mathbf{F}}(\pi_{\infty}^{\star})=\lambda\mathcal{L}_{\mathbf{F}}(\pi_{\infty}^{\star})-2\lambda\sum_{(s,w)\in[K_{s}]\times[K_{w}]}\big\langle\langle\xi_{sw},\pi^{\star}_{\infty}\rangle\xi_{sw},\pi_{\infty}^{\star}\big\rangle,$
because, for any $(s,w)\in\{0,1\}^{2}$, $\big\langle\langle\xi_{sw},\pi^{\star}_{\infty}\rangle\xi_{sw},\pi_{\infty}^{\star}\big\rangle=\big\langle\xi_{sw},\pi_{\infty}^{\star}\big\rangle^{2}$, hence
$\mathcal{L}_{\mathbf{F}}(\pi_{\infty}^{\star})=\sum_{(s,w)\in[K_{s}]\times[K_{w}]}\big\langle\langle\xi_{sw},\pi^{\star}_{\infty}\rangle\xi_{sw},\pi_{\infty}^{\star}\big\rangle.$
This, in turn, reduces the proof of (16) to showing
$\mathbb{E}[|\lambda\mathcal{L}_{\mathbf{F}}(\hat{\pi}_{n}^{\star})-\lambda\mathcal{L}_{\mathbf{F}}(\pi_{\infty}^{\star})|]\leq\mathcal{O}(1/\sqrt{n})$ (17)
and
$\mathbb{E}\Big[2\lambda\big|\sum_{(s,w)\in[K_{s}]\times[K_{w}]}\big\langle\langle\xi_{sw},\pi^{\star}_{\infty}\rangle\xi_{sw},\pi_{\infty}^{\star}-\hat{\pi}_{n}^{\star}\big\rangle\big|\Big]\leq\mathcal{O}(1/\sqrt{n}).$ (18)
To obtain these two inequalities and complete the proof, it suffices to note that the conclusion of Lemma B.2 allows us to apply Lemma˜B.4 (sub-Gaussian concentration, which is stronger, readily implies the expectation bound used here), which yields the desired inequality (17). To obtain inequality (18) just note that, for $s,w\in[K_{s}]\times[K_{w}]$,
$\langle\xi_{sw},\pi^{\star}_{\infty}\rangle^{2}-\langle\xi_{sw},\hat{\pi}_{n}^{\star}\rangle^{2}=\langle\xi_{sw},\pi^{\star}_{\infty}-\hat{\pi}_{n}^{\star}\rangle\langle\xi_{sw},\pi^{\star}_{\infty}+\hat{\pi}_{n}^{\star}\rangle$
which implies that
$\big|\langle\xi_{sw},\pi^{\star}_{\infty}\rangle^{2}-\langle\xi_{sw},\hat{\pi}_{n}^{\star}\rangle^{2}\big|\leq 2\|\xi_{sw}\|_{\infty}\big|\langle\xi_{sw},\pi^{\star}_{\infty}-\hat{\pi}_{n}^{\star}\rangle\big|;$
and the bound in expectation is once more a matter of invoking Lemma˜B.4. ∎
B.2.1 Technical Lemmas
Lemma B.2.
Let $\pi^{\star}_{\infty}$ be a minimizer of
$\min_{\begin{subarray}{c}\pi_{1}=\mu\\ \pi_{2}=\eta\end{subarray}}F(\pi)\vcentcolon=\langle c,\pi\rangle+\varepsilon\mathbf{KL}\big(\pi||\mu\otimes\eta\big)+\lambda\mathcal{L}_{\mathbf{F}}(\pi):=m_{\infty}^{\star}.$ (19)
Then $\pi^{\star}_{\infty}$ also minimizes
$\min_{\begin{subarray}{c}\pi_{1}=\mu\\ \pi_{2}=\eta\end{subarray}}F_{L}(\pi)\vcentcolon=\langle c,\pi\rangle+\varepsilon\mathbf{KL}(\pi||\mu\otimes\eta)+\lambda\mathcal{L}_{\mathbf{F}}(\pi^{\star}_{\infty})+\big\langle 2\lambda\sum_{s,w\in[K_{s}]\times[K_{w}]}\langle\xi_{sw},\pi^{\star}_{\infty}\rangle\xi_{sw},\pi-\pi^{\star}_{\infty}\big\rangle$ (20)
$=$ $\min_{\begin{subarray}{c}\pi_{1}=\mu\\ \pi_{2}=\eta\end{subarray}}\langle c+2\lambda\sum_{s,w\in[K_{s}]\times[K_{w}]}\langle\xi_{sw},\pi_{\infty}^{\star}\rangle\xi_{sw},\pi\rangle+\varepsilon\mathbf{KL}(\pi||\mu\otimes\eta)-\lambda\mathcal{L}_{\mathbf{F}}(\pi_{\infty}^{\star})$
$=$ $\widehat{m}_{\infty}^{\star}-\lambda\mathcal{L}_{\mathbf{F}}(\pi^{\star}_{\infty}).$
and, consequently, $\widehat{m}_{\infty}^{\star}-\lambda\mathcal{L}_{\mathbf{F}}(\pi^{\star}_{\infty})=m_{\infty}^{\star}.$
Proof.
The proof is inspired by the one from [rakotomamonjy2015generalized, Proposition 2.1]. Assume $\pi_{\infty}^{\star}$ minimizes Eq.˜19. The presence of entropic regularization implies that $\pi_{\infty}^{\star}\ll\mu\otimes\eta$. Let
$g_{L}=2\sum_{s,w\in[K_{s}]\times[K_{w}]}\langle\xi_{sw},\pi_{\infty}^{\star}\rangle\xi_{sw}.$
We already know that for any $\pi\in\Pi(\mu,\eta)$
$\mathcal{L}_{\mathbf{F}}(\pi)\geq\mathcal{L}_{\mathbf{F}}(\pi_{\infty}^{\star})+\langle g_{L},\pi-\pi_{\infty}^{\star}\rangle.$
Define
$g_{\text{KL}}=\log\left(\frac{d\pi_{\infty}^{\star}}{d(\mu\otimes\eta)}\right),$
we prove, for any $\pi\in\Pi(\mu,\eta)$,
$\mathbf{KL}(\pi||\mu\otimes\eta)\geq\mathbf{KL}(\pi_{\infty}^{\star}||\mu\otimes\eta)+\langle g_{\text{KL}},\pi-\pi_{\infty}^{\star}\rangle,$
with convex analysis [^4] If $\pi$ is singular with respect to $\mu\otimes\eta$, the inequality is trivial. Assume $\pi\ll\mu\otimes\eta$, we define for $r\in[0,1]$, $\pi_{r}=(1-r)\pi_{\infty}^{\star}+r\pi\in\Pi(\mu,\eta)$ and $\phi(r)=\mathbf{KL}(\pi_{r}||\mu\otimes\eta)$. By convexity of $\phi$, $\phi(r)\geq\phi(0)+r\lim_{h\to 0^{+}}\frac{\phi(h)-\phi(0)}{h}$. We compute the previous directional limit using the monotone convergence theorem as in CSZISAR_SUBDIFF., and thus
$F(\pi)\geq F(\pi_{\infty}^{\star})+\langle c+\lambda g_{L}+\varepsilon g_{\text{KL}},\pi-\pi_{\infty}^{\star}\rangle.$
Consequently, for any $\pi\in\Pi(\mu,\eta)$ such that $\pi\ll\mu\otimes\eta$ [^5] If instead $\pi$ is singular with respect to $\mu\otimes\eta$, then $F(\pi)=F_{L}(\pi)=\infty$, so such measures are trivially excluded from being minimizers of either functional.,
$\langle c+\lambda g_{L}+\varepsilon g_{\text{KL}},\pi-\pi_{\infty}^{\star}\rangle\geq 0,$
otherwise, for $r\in[0,1]$, $\pi_{r}=(1-r)\pi_{\infty}^{\star}+r\pi\in\Pi(\mu,\eta)\cap\{\eta;\eta\ll\mu\otimes\eta\}$, we could find $r^{\star}\in(0,1]$, such that $F(\pi_{r^{\star}})<F(\pi_{\infty}^{\star})$. In fact, by contradiction, assume $\langle c+\lambda g_{L}+\varepsilon g_{\text{KL}},\pi-\pi_{\infty}^{\star}\rangle<0$, as $\lim_{r\to 0^{+}}\frac{F(\pi_{r})-F(\pi_{\infty}^{\star})}{r}=\langle c+\lambda g_{L}+\varepsilon g_{\text{KL}},\pi-\pi_{\infty}^{\star}\rangle$, we have, by continuity, the existence of a sufficiently small $r^{\star}\in(0,1]$ such that $F(\pi_{r^{\star}})<F(\pi_{\infty}^{\star})$. Moreover, as $F_{L}(\pi_{\infty}^{\star})=F(\pi_{\infty}^{\star})$, for any $\pi\in\Pi(\mu,\eta)$
$F_{L}(\pi)\geq F_{L}(\pi_{\infty}^{\star})+\langle c+\lambda g_{L}+\varepsilon g_{\text{KL}},\pi-\pi_{\infty}^{\star}\rangle\geq F_{L}(\pi_{\infty}^{\star}).$
Hence, $\pi_{\infty}^{\star}$ is also a minimizer of Eq.˜20. ∎
B.3 Proof of Section˜4.2
We begin by showing the following Lemma, which mimics the arguments of rigollet2022sample.
Lemma B.3.
Let $f_{\star},g_{\star}$ be dual potentials such that $\int g_{\star}(y)d\eta(y)=0$. Then we have
$\lVert f_{\star}\rVert_{\infty}\leq\lVert c_{\theta}\rVert_{\infty}$
$\lVert g_{\star}\rVert_{\infty}\leq\lVert c_{\theta}\rVert_{\infty}.$
Proof.
Let $f_{\star},g_{\star}$ be two dual potentials. Since
$\int e^{-\frac{1}{\varepsilon}\big(c_{\theta}(x,y)-f_{\star}(x)-g_{\star}(y)\big)}d\eta(y)=1,$
we have, by using the fact that $\eta$ and $\mu$ have bounded support, that
$1=\int e^{-\frac{1}{\varepsilon}\big(c_{\theta}(x,y)-f_{\star}(x)-g_{\star}(y)\big)}d\eta(y)\geq e^{-\frac{1}{\varepsilon}\big(\max\limits_{x,y}c_{\theta}(x,y)-f_{\star}(x)\big)}\int e^{\frac{1}{\varepsilon}g_{\star}(y)}d\eta(y).$
By Jensen’s inequality,
$\int e^{\frac{1}{\varepsilon}g_{\star}(y)}d\eta(y)\geq e^{\frac{1}{\varepsilon}\int g_{\star}(y)d\eta(y)}=1$
using the convention $\int g_{\star}(y)d\eta(y)=0$. We are thus left with
$e^{-\frac{1}{\varepsilon}\big(\max\limits_{x,y}c_{\theta}(x,y)-f_{\star}(x)\big)}\leq 1$
which can be rearranged to yield
$f_{\star}(x)\leq\lVert c_{\theta}\rVert_{\infty}$
$\mu$-almost everywhere where $\lVert c_{\theta}\rVert_{\infty}:=\max\limits_{x,y}c(x,y)$. We may now using this inequality to write that
$f_{\star}(x)+g_{\star}(y)-c_{\theta}(x,y)\leq f_{\star}(x)+g_{\star}(y)\leq g_{\star}(y)+\lVert c_{\theta}\rVert_{\infty}$
since $c_{\theta}(x,y)\geq 0$. This gives
$e^{-\frac{1}{\varepsilon}\big(c_{\theta}(x,y)-f_{\star}(x)-g_{\star}(y)\big)}\geq e^{-\frac{1}{\varepsilon}\big(g_{\star}(y)+\lVert c_{\theta}\rVert_{\infty}\big)}.$
Using again the fact that the left-hand side of this inequality is equal to $1$, we get
$g_{\star}(y)\geq-\lVert c_{\theta}\rVert_{\infty}.$
Leveraging the same argument than rigollet2022sample, we have
$\int f_{\star}(x)d\mu(x)\geq 0,$
which in turn allows us to conclude by symmetrical arguments and to get
$\lVert f_{\star}\rVert_{\infty}\leq\lVert c_{\theta}\rVert_{\infty}$
$\lVert g_{\star}\rVert_{\infty}\leq\lVert c_{\theta}\rVert_{\infty}.$
Using this Lemma, we get the following result which extends Theorem 6 from rigollet2022sample to parametrized costs.
Lemma B.4.
Let $\varphi\in L^{\infty}(\mu\otimes\eta)$. For all $t>0$, with probability at least $1-18e^{-t^{2}}$, we have
$\big\lvert\int\varphi d(\pi_{n}-\pi_{\star})\big\lvert\lesssim\frac{\exp\big(5R_{\Theta}\varepsilon^{-1}\big)\lVert\varphi\rVert_{L^{\infty}(\mu\otimes\eta)}\cdot t}{\sqrt{n}}.$
Proof.
We may now write
$p_{\star}(x,y)=e^{-\frac{\lVert c_{\theta}\rVert_{\infty}}{\varepsilon}\big(\lVert c_{\theta}\rVert_{\infty}^{-1}c_{\theta}(x,y)-\lVert c_{\theta}\rVert_{\infty}^{-1}f_{\star}(x)-\lVert c_{\theta}\rVert_{\infty}^{-1}g_{\star}(y)\big)}.$
This rescaling allows us to get back to the setup of rigollet2022sample, who derive from their assumptions that
$c_{\theta}(x,y)\leq 1,\quad f_{\star}(x)\leq 1\quad\textnormal{and}\quad g_{\star}(x)\leq 1,$
and use these bounds in the rest of their proof. Hence, by rescaling the entropic penalization strength as $\varepsilon^{-1}\lVert c_{\theta}\rVert_{\infty}$, we obtain
$e^{-\frac{5\lVert c_{\theta}\rVert_{\infty}}{\varepsilon}}\leq p_{\star}(x,y)\leq e^{\frac{5\lVert c_{\theta}\rVert_{\infty}}{\varepsilon}}$
from the proof of Proposition 10 of rigollet2022sample. We finish the proof similarly as Theorem 6 from rigollet2022sample. ∎
We are now ready to proceed with the proof of Theorem 4.2.
Proof.
For simplicity, we consider the case where the fairness penalty is independent of the cost function. Using the identity $\big\lvert a^{2}-b^{2}\big\lvert\leq\big\lvert(a+b)\big\lvert\big\lvert(a-b)\big\lvert$, we first have that for all $\theta\in\Theta$ that
$\big\lvert\mathcal{L}_{\mathbf{F}}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})\big)-\mathcal{L}_{\mathbf{F}}\big(\pi^{\star}_{\epsilon}(c_{\theta})\big)\big\lvert\lesssim\sum\limits_{s=1}^{K_{s}}\sum\limits_{w=1}^{K_{w}}\big\lvert\textnormal{Tr}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})^{\top}\mathbf{B}_{sw}\big)-\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi^{\star}_{\epsilon}(c_{\theta})(x,u,y,t)\big\lvert$
since
$\sum_{s=1}^{K_{s}}\sum_{w=1}^{K_{w}}\big\lvert\textnormal{Tr}\big(\mathbf{\Pi}_{\epsilon}^{\top}(c_{\theta})\mathbf{B}_{sw}\big)+\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi^{\star}_{\epsilon}(c_{\theta})(x,u,y,t)+2\mathbf{F}_{sw}\big\lvert$
can be upper-bounded by a constant. Now, for any $(s,w)\in[1,K_{s}]\times[1,K_{w}]$ we denote
$\varphi_{sw}~=~\xi_{sw}+~\mathbf{F}_{sw}$
where $\xi_{sw}$ is defined by Equation (10). We observe that $\varphi_{sw}\in L^{\infty}(\mu\otimes\eta)$. Thus, by using Lemma˜B.4 we have with probability at least $1-18e^{-t^{2}}$
$\big\lvert\textnormal{Tr}\big(\mathbf{\Pi}_{\epsilon}^{\top}(c_{\theta})\mathbf{B}_{sw}\big)-\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi^{\star}_{\epsilon}(c_{\theta})(x,u,y,t)\big\lvert$ $=\big\lvert\int\varphi_{sw}d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\lvert$
$\lesssim\frac{\exp(5R_{\Theta})\epsilon^{-1}\cdot t}{\sqrt{n}}$ (21)
The random variable $X_{sw}:=\big\lvert\int\varphi_{sw}d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert$ being non-negative, we have
$\mathbb{E}\big[X_{sw}\big]=\int_{0}^{\infty}\mathbb{P}[X_{sw}>u]du.$
Letting $C_{sw}$ the constant hidden by $\lesssim$ in (21), with the change of variable $u=\frac{C_{sw}\exp(5R_{\Theta})\epsilon^{-1}\cdot t}{\sqrt{n}}$, we get
$\mathbb{E}\big[X_{sw}\big]$ $=\frac{C_{sw}\exp(5R_{\Theta})\epsilon^{-1}\cdot}{\sqrt{n}}\int_{0}^{\infty}\mathbb{P}\big[X_{sw}>\frac{C_{sw}\exp(5R_{\Theta})\epsilon^{-1}\cdot t}{\sqrt{n}}\big]dt$
$\leq\frac{C_{sw}\exp(5R_{\Theta})\epsilon^{-1}\cdot}{\sqrt{n}}\int_{0}^{\infty}\min(1,18e^{-t^{2}})dt$
Let $t_{0}=\sqrt{\log(18)}$. It follows
$\mathbb{E}\big[X_{sw}\big]$ $\leq\frac{C_{sw}\exp(5R_{\Theta})\epsilon^{-1}\cdot}{\sqrt{n}}\big[t_{0}+\int_{t_{0}}^{\infty}18e^{-t^{2}}dt\big]$
$\leq\frac{C_{sw}\exp(5R_{\Theta})\epsilon^{-1}\cdot}{\sqrt{n}}\big[t_{0}+\frac{e^{-t_{0}^{2}}}{2t_{0}}\big]$
$\leq\frac{\tilde{C_{sw}}\exp(5R_{\Theta})\epsilon^{-1}\cdot}{\sqrt{n}}$
where the second inequality comes from an integration by parts. Summing over $s$ and $w$ yields for any $\theta\in\Theta$
$\mathbb{E}[\lvert\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\mathcal{L}_{\mathbf{F}}(\pi^{*}_{\epsilon}(c_{\theta}))\rvert]\lesssim\frac{\exp(5R_{\Theta})\epsilon^{-1}}{\sqrt{n}}$
As a consequence
$\sup_{\theta\in\Theta}\mathbb{E}[\lvert\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\mathcal{L}_{\mathbf{F}}(\pi^{*}_{\epsilon}(c_{\theta}))\rvert]\lesssim\frac{\exp(5R_{\Theta})\epsilon^{-1}}{\sqrt{n}}$
Appendix C Computational Details
C.1 Implementation Details
C.1.1 Fair Sinkhorn
We implement the FairSinkhorn algorithm from scratch using Sinkhorn-Knopp matrix scaling as proposed in cuturi2013sinkhorn.
C.1.2 Penalized Optimal Transport
We leverage the Python OT (POT) library [flamary2021pot] throughout our experiments. To solve the penalized problem, we leverage the generic conditional gradient algorithm implemented in this package. We use the following parameters and refer to the package documentations for greater details.
Parameter Value
G0 Initialized via a solution of the entropic OT problem
numIterMax $10$
numInnerItermax $200$
stopThr $10^{-9}$
stopThr2 $10^{-9}$
Table 1: Parameters of the gcg algorithm used.
The algorithm is used in conjunction with line_search_armijo provided in Python OT.
C.1.3 Cost Learning
We implement bilevel cost learning in Pytorch. The inner optimal transport problem is solved using the ot.sinkhorn method from POT, with a warm-started initialization of the dual variables, a stoping threshold of $10^{-6}$ and a maximum of $1000$ iterations. The used solver is sinkhorn if the entropic penalty is greater than $1$, otherwise we use sinkhorn_log. The Jacobian of the optimal transport plan is estimated by iterative differentiation [Pauwels2023]. We take a single optimizer step after each solve of the inner problem. We found Adam [kingma2014adam] to provide the best and most stable results over runs. We use its default parameters, except for the learning rate. In all experiments, we set the penalty $\mathscr{D}(\cdot,\cdot)$ to be the squared Frobenius norm between the original and parametrized cost matrix, divided by the maximum entry of the euclidean cost matrix.
C.2 Experimental Details
To generate results in Figure 3, we generate $n_{x}=n_{y}=20$ samples for each subgroup of the $2$ subgroups in each distribution. For cost learning, we use a penalty grid of $\texttt{logspace}(-3,2,30)$ and set $\varepsilon=2$. The learning rate is set to $0.1$ for the Mahalanobis and ResNet cost models, and to $0.05$ for the MLP cost model. The MLP model has $3$ hidden layers and a output dimension of $8$. The dimension of the intermediate states is set to $4$. The same architectural parameters are used for the ResNet model. We use a SiLU activation for both models. All cost learning models are trained for maximum $300$ iterations of the outer problem. For the penalized OT results, we use a penalty grid of $\texttt{logspace}(0,3,20)$ and set $\varepsilon=2$.
To generate results in Figure 5, we generate $n_{x}=n_{y}=20$ samples for each subgroup of the $3$ subgroups in each distribution. For cost learning, we use a penalty grid of $\texttt{logspace}(-3,2,30)$ and set $\varepsilon=2$. The learning rate is set to $0.1$ for all cost learning models. The MLP model has $3$ hidden layers and a output dimension of $8$. The dimension of the intermediate states is set to $16$. The same architectural parameters are used for the ResNet model. We use a SiLU activation for both models. All cost learning models are trained for maximum $300$ iterations of the outer problem. For the penalized OT results, we use a penalty grid of $\texttt{logspace}(-3,3,20)$ and set $\varepsilon=2$.
To generate results in Figure 7, we set a fixed penalty of $0.1$, a fixed entropic penalty of $\varepsilon=2$ and a fixed learning rate of $0.01$. The Mahalanobis and ResNet cost model are used with identical parameters to Figure 3. We learn a cost model on the same dataset than for Figure 3 but with $1000$ samples per subgroup. We then evaluate the fairness of each cost model on small subsamples of $50$ samples per subgroup. The results are computed for $10$ runs. For penalized OT, we tune the penalty value to $15$ to obtain similar levels of fairness, and use the same entropic penalty $\varepsilon=2$.

### v2

@@
 Focusing on Optimal Transport (OT), we introduce a novel notion of group fairness requiring that the probability of matching two individuals from any two given groups in the OT plan satisfies a predefined target.
-We first propose FairSinkhorn, a modified Sinkhorn algorithm to compute perfectly fair transport plans efficiently.
+We first propose a modified Sinkhorn algorithm to compute perfectly fair transport plans efficiently.
 Since exact fairness can significantly degrade matching quality in practice, we then develop two relaxation strategies.
-The first one involves solving a penalised OT problem, for which we derive novel finite-sample complexity guarantees.
-This result is of independent interest as it can be generalized to arbitrary convex penalties.
-Our second strategy leverages bilevel optimization to learn a ground cost that induces a fair OT solution, and we establish a bound guaranteeing that the learned cost yields fair matchings on unseen data.
-Finally, we present empirical results that illustrate the trade-offs between fairness and performance.
+The first one involves solving a penalized OT problem, for which we derive novel finite-sample complexity guarantees.
+Our second strategy leverages bilevel optimization to learn a ground cost that induces a fair OT solution, and we establish a bound on the deviation of fairness when matching unseen data.
+Finally, we present empirical results illustrating the performance of our approaches and the trade-off between fairness and transport cost.
 1 Introduction
-Algorithmic matching mechanisms play an increasing role in modern societies, handling the distribution of rare goods by connecting individuals or firms through algorithmic decisions rather than through price-driven markets.
-Instead of choosing an exchange partner through a decentralized market in which prices make matches, participants are allocated one or multiple partners through a centralized clearing house that takes into account their individual characteristics and preferences.
+Algorithmic matching mechanisms play an increasing role in modern societies, handling the distribution of rare goods by centrally connecting individuals or firms based on their characteristics and preferences rather than through price-driven markets.
 Examples of such mechanisms include online job recommendations, college admissions systems, dating and ride-hailing apps, and kidney allocation circuits.
@@
 For instance, the growing importance of dating apps in contemporary matrimonial dynamics has raised concerns over their role in fostering social and racial marital homogamy [zheng2018fairness, jia2018online, zhao2024leveraging], a phenomenon often highlighted as a key driver of inequality.
-Similarly, school admission systems have been criticized for excessively matching students from privileged backgrounds with elite institutions regardless of their academic potential [hiller2012choix, fack2014impact, simioni2022comment]—causing underprivileged students to lose the benefits of elite higher education.
-More generally,
-these concerns arise because matching decisions overly depend on individual characteristics that are highly correlated with attributes defining social groups, such as ethnicity or socioeconomic background.
+Similarly, school admission systems have been criticized for excessively matching students from privileged backgrounds with elite institutions regardless of their academic potential [hiller2012choix, fack2014impact, simioni2022comment], causing underprivileged students to lose the benefits of elite higher education.
+More generally, these concerns arise because matching decisions overly depend on individual characteristics that are highly correlated with attributes defining social groups, such as ethnicity or socioeconomic background.
 Such dependence may lead to highly homogeneous pairings, which may conflict with political, philosophical, and legal principles that promote diversity and equality of opportunity.
@@
 In supervised learning, where an algorithm predicts an unknown outcome $Y$ from features $X$, one example of group fairness is demographic parity [calders2009building], which requires that the prediction $f(X)$ be independent of a protected attribute $S$ (such as gender or age).
-In practice, this means that predictions for individuals—such as credit scores—should not systematically depend on sensitive attributes.
 In unsupervised learning, one approach is equality of representation error, which requires that the reconstruction error of a learned latent representation be independent of the sensitive attribute $S$ [samadi2018price].
@@
 Extending these fairness notions to matching problems is, however, nontrivial, as matching cannot be directly framed as either a supervised or unsupervised learning task.
-Most prior work has focused on individual fairness (ensuring that similar individuals receive similar matches) or on problem-specific notions, such as guaranteeing a minimal participation level in settings where individuals can opt out of the matching mechanism.
+Most prior work has focused on individual fairness, ensuring that similar individuals receive similar matches [devic2023fairness], or on problem-specific notions, such as guaranteeing a minimal participation level in settings where individuals can opt out [ashlagi2014free].
 We discuss this literature in more details in the related work section.
+Figure 1: Illustration of optimal transport under group fairness constraints.
+Left: Due to the correlation between group attributes ($S$ and $W$) and features, vanilla optimal transport produces transport plans that concentrate most of the mass between matching groups.
+Center: Our novel notion of fairness is based on fairness targets, defined as the desired group-level matching probabilities specified by a central planner.
+Right: We propose two approaches for solving this problem: a fairness-penalized optimal transport method, and a cost-learning approach that transforms the input space so that the resulting optimal transport plan aligns as closely as possible with the fairness targets.
 Contributions.
 In this work, we study group fairness in matching using the framework of optimal transport (OT).
-OT has been a long standing modeling tool in economics and social sciences since the pioneering work of Kantorovitch, eventually earning him the Nobel Memorial Prize in Economic Sciences in 1975 [galichon2018optimal, galichon2021unreasonable].
-It has gained traction in the last decade as a way of studying complex matching mechanisms, finding applications in the economics of international trade [galichon2021unreasonable, mastrandrea2025optimal], logistics and urban network studies [leite2022urban], demand estimation and pricing [qu2025sinkhorn], decision theory, as well as labor and migration economics [carlier2023sista, hazard2025whom].
-We introduce a novel notion of fairness for optimal transport that allows a central planner—such as a government or regulatory body—to specify the probability that individuals from one subgroup are matched with individuals from another subgroup.
-This flexible definition can enforce a variety of constraints, including limited homogamy (reducing the likelihood that individuals with similar sensitive attributes are matched), matching quotas (ensuring that $p$ percent of a minority subgroup is matched to another group), or broader goals of social diversity.
-Specifically, our contributions are as follows.
+Since the pioneering work of Kantorovitch, OT has been a long standing modeling tool in economics and social sciences [galichon2018optimal] and has gained traction in the last decade as a way of studying complex matching mechanisms in economics [galichon2021unreasonable, mastrandrea2025optimal], logistics and urban network studies [leite2022urban], demand estimation and pricing [qu2025sinkhorn], decision theory, as well as labor and migration economics [carlier2023sista, hazard2025whom].
+We introduce a novel notion of fairness for optimal transport that allows a central planner—such as a government or regulatory body—to specify the probability that individuals from one group are matched with individuals from another group.
+This flexible definition can enforce a variety of constraints, including limited homogamy (reducing the likelihood that individuals with the same sensitive attributes are matched), matching quotas (ensuring that a certain percentage of a minority group receives desirable opportunities), or broader goals of social diversity.
+We summarize our contributions below, with an overview provided in Figure 1.
   1. 1.
-    We first provide a theoretical framework for fair optimal transport, introducing a formal and novel definition.
-Building on this framework, we present the FairSinkhorn algorithm for computing optimal transport plans under exact fairness constraints.
+    We introduce a novel fair optimal transport problem, and present the FairSinkhorn algorithm for computing OT plans under exact fairness constraints.
   2. 2.
-    We design two relaxed approaches.
-The first one penalizes the optimal transport objective with a convex penalty that promotes fair transport plans.
-This formulation yields a convex optimization problem in which the utility–fairness trade-off can be easily controlled by adjusting the fairness penalty parameter.
-The second approach consists in learning a cost function such that the resulting transport plan is fair.
-It has the advantage of enabling to easily match new samples.
+    We propose two relaxed approaches.
+The first adds a penalty to the optimal transport objective to encourage fair plans, yielding a convex optimization problem where the cost–fairness trade-off is easily controlled via the penalty parameter.
+The second learns a cost function such that the resulting transport plan is fair, allowing straightforward matching of new samples.
   3. 3.
     For the first approach, we provide a novel sample complexity bound for penalized entropic optimal transport, which we believe to be of independent interest.
-For the cost learning approach, we provide a bound on the expected fairness deviation when using finite sample to compute the transport plan with the learned cost function.
+For the cost learning approach, we bound the expected fairness deviation between finite samples and the distribution, holding for any cost in a given parameterized family.
   4. 4.
-    Lastly, we investigate the empirical performance of the proposed approaches through a series of numerical experiments.
+    We investigate the empirical performance of the proposed approaches through numerical experiments.
 Our code will be made public upon publication. Throughout this article, all proofs are deferred to the appendix.
 2 Related Work
-Fairness in Matching Mechanisms.
-Since the pioneering work of gale1962college on the stable marriage problem, a series of works have provided theoretical studies of the fairness in matching mechanisms.
-A first line of research focuses on the stability of GS-like algorithms under constraints akin to individual fairness, ensuring that similar individuals receive similar outcomes by the decision-making process.
-For instance, karni2021fairness analyze the Gale-Shapley (GS) algorithm through the lens of preference informed individual fairness.
-This criterion allows for small deviations from individual fairness, which requires that similar individuals should be treated similarly.
-They devise a fair variant of the GS algorithm that ensure both approximately fair and stable matchings.
-Similarly, devic2023fairness analyze merit-based fairness of one-to-one matching mechanisms, a notion close to individual fairness, under partial uncertainty over individual’s merit.
-A second line of research introduces metrics of fairness closer to group fairness, with a specific focus on kidney paired donation programs, a singular matching problem subject to domain specific constraints — in these programs, donors and patients must be matched by forming cycles due to legal constraints, which leads to a reformulation of the matching problem as a constrained integer programming problem.
-In their seminal work, ashlagi2014free unveil an unfair market collapse phenomenon in kidney pair donation programs: instead of introducing all patients to the matching system, hospitals internally operate easy matches i.e.
-do not participate fully in the exchange mechanism.
-This leads to a loss in transplantation opportunities for highly-sensitized patients that are hard to match.
-dickerson2014price and following work by zhang2025fairness also consider the case of kidney paired donations programs.
-Their definition of fairness enforces a minimal amount of program participation for the group of highly-sensitized patients.
-More recently, lobo2025fair consider a combinatorial multimatching problem under uncertainty: in their setup, the goal is to match agents and resources without exact knowledge of their pairwise valuations.
-Their definition of fairness relies on the average utility within predefined groups, which is closest to our definition among previous work.
-Our contribution differentiates from these works in two aspects.
-First, we do not consider iterative matching mechanisms such as the GS algorithm: our work focuses on matching through (entropic) optimal transport.
-Secondly, our definition of fairness is different from previously introduced definitions, and highly modular to accommodate fairness preferences of a central planner.
-Instead of requiring individual fairness, building on individual rankings over possible assignements or participation constraints, we consider a matching problem between two populations (f.i.
-students and universities) further structured by a sensitive attribute (f.i.
-privileged and unprivileged students, and elite and non-elite universities).
-We enforce matching quotas given by a central planner, that specify for each subgroup the proportion of individuals that should be matched to subgroups in the second population.
+Fairness in matching algorithms.
+Fairness has been studied across a variety of matching problems.
+A first line of work investigates individual-level fairness in Gale–Shapley–type matching mechanisms [gale1962college], requiring that similar individuals receive similar outcomes [karni2021fairness, devic2023fairness].
+A second line of work focuses on group-level fairness, primarily in the specific case of kidney paired donation (KPD) programs [ashlagi2014free, dickerson2014price, zhang2025fairness, lobo2025fair].
+Our work departs from this literature in two key ways.
+First, we do not consider iterative agent-level mechanisms (in which agents are sequentially paired until the match is stable) but instead model matching through (entropic) optimal transport.
+Second, rather than individual rankings, we assume that both populations are structured by sensitive attributes and impose explicit inter-group mass constraints specifying how much probability mass should flow between subgroups.
+Another related line of work studies fairness-constrained bipartite matching on graphs $\big(\mathcal{U}\cup\mathcal{V},\mathcal{E}\big)$, where edges encode feasible matches [godsil1981matchings, zdeborova2006number, noiry2021online].
+Recent work in this setting includes [castera2024price], who partition one side of the graph ($\mathcal{U}$ or $\mathcal{V})$ into groups and limit the number of vertices that can be matched from each group, and [panda2024individual], who further impose individual fairness constraints.
+Our work is fundamentally different from theirs since $(i)$ individuals are represented by features rather than graph positions; $(ii)$ we allow mass splitting rather than integral matchings; and $(iii)$ our formulation is symmetric in the two populations.
 Fairness and optimal transport.
-Recent work has drawn many fruitful connection between fairness and optimal transport, primarily by using OT as a tool to obtain or characterize fair learning algorithms [gordaliza2019obtaining, gouic2020projection, chzhen2020fair, DBLP:conf/aaai/ChiappaJSPJA20, DBLP:conf/nips/BuylB22, hu2023fairness, DBLP:conf/icml/XianY023, chowdhary2024fairness, xiong2024fair, divol2024demographic].
-A central insight of this line of work is that learning an optimal fair predictor can be formulated as a Wasserstein barycenter problem [gouic2020projection, chzhen2020fair, divol2024demographic].
-This perspective is orthogonal to ours: OT serves as an auxiliary mechanism for analyzing and enforcing fairness in downstream prediction tasks, whereas our goal is to study and enforce fairness of the optimal transport itself.
-Closer in spirit to our work is the recent article by nguyen2024towards, which imposes fairness constraints directly on a Wasserstein barycenter by requiring approximate equality of sliced-Wasserstein distances SW [nadjahi2020statistical] to multiple marginals $\mu_{1},\dots,\mu_{K}$:
-$\min\limits_{\mu}\sum\limits_{k=1}^{K}\textnormal{{SW}}(\mu,\mu_{k})\textnormal{ s.t }\sum\limits_{i=1}^{K}\sum\limits_{j=1}^{K}\big\lvert\textnormal{{SW}}(\mu,\mu_{i})-\textnormal{{SW}}(\mu,\mu_{j})\big\lvert\leq\varepsilon.$
-Their approach, however, does not consider sensitive attributes or demographic groups, focusing instead on balancing distances across marginals.
+A growing literature uses optimal transport to obtain or characterize fair learning algorithms [gordaliza2019obtaining, gouic2020projection, chzhen2020fair, DBLP:conf/aaai/ChiappaJSPJA20, DBLP:conf/nips/BuylB22, hu2023fairness, DBLP:conf/icml/XianY023, xiong2024fair, divol2024demographic].
+A central insight is that fair prediction can often be formulated as a Wasserstein barycenter or projection problem [gouic2020projection, chzhen2020fair, divol2024demographic].
+In this line of work, OT is a tool to enforce fairness in downstream prediction tasks.
+Our perspective is orthogonal: we study the fairness of the transport plan itself.
+Closest to our problem is nguyen2024towards, who impose fairness on a Wasserstein barycenter problem by enforcing approximate equality of its sliced-Wasserstein distances to the input measures.
 In contrast, we study the problem of computing transport plans that satisfy explicit mass constraints between groups defined by sensitive attributes, a setting that, to the best of our knowledge, has not been previously considered.
 Constrained and statistical optimal transport.
-Our work borrows from recent developments in constrained optimal transport, which seeks to enforce structural properties on the transport plan [courty2016optimal, blondel2018smooth, paty2019subspace, liu2022sparsity]; for instance, korman2015optimal analyze a variant of optimal transport in which the amount of mass that can be transported between two units is upper bounded.
-Specifically, we build on rakotomamonjy2015generalized, genevay2019sample and rigollet2022sample to derive finite sample guarantees for a penalized optimal transport problem, and to bound the deviations of fairness with learned cost functions.
-3 Formalizing Group Fairness in Optimal Transport
+Our approach builds on constrained optimal transport, which enforces structural properties of transport plans [courty2016optimal, blondel2018smooth, paty2019subspace, liu2022sparsity], including bounds on transported mass [korman2015optimal].
+We build on and extend the theory of statistical and penalized optimal transport [rakotomamonjy2015generalized, genevay2019sample, rigollet2022sample] to derive finite-sample guarantees and to control deviations from the prescribed fairness constraints.
+We provide an extended related work review in Appendix A.
+3 Formalizing Group Fairness in OT
 3.1 Optimal Transport
-Entropic Optimal Transport.
-We consider the entropic regularized optimal transport problem [peyre2019computational, genevay2019sample, feydy2019interpolating, pooladian2021entropic, keriven2022entropic] between two probability distributions $\mu$ and $\eta$ with ground cost $c$:
-$\mathbf{W}_{c,\varepsilon}(\mu,\eta):=\min\limits_{\pi\in\Pi(\mu,\eta)}\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\,d\pi(x,y)+\varepsilon\mathbf{KL}\big(\pi|\mu\otimes\eta\big),$
+We consider the entropic optimal transport problem [peyre2019computational, genevay2019sample, feydy2019interpolating, pooladian2021entropic, keriven2022entropic] between two probability distributions $\mu$ and $\eta$ with ground cost $c$:
+$\min\limits_{\pi\in\Pi(\mu,\eta)}\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\,d\pi(x,y)+\varepsilon\mathbf{KL}\big(\pi|\mu\otimes\eta\big),$
 where the minimum is taken over couplings $\Pi(\mu,\eta)$ of $\mu$ and $\eta$ and where $\mathbf{KL}$ denotes the Kullback-Leibler (KL) divergence [CSZISAR_SUBDIFF].
-We focus on the entropic formulation because the regularization ensures uniqueness and smoothness of the optimal coupling, which is crucial for efficient computation and gradient-based optimization [peyre2019computational].
-With a slight abuse of notation, we will often omit the explicit dependence on $\mu$ and $\eta$ when clear from context, writing, for example, $\mathbf{W}_{c,\varepsilon}$ instead of $\mathbf{W}_{c,\varepsilon}(\mu,\eta)$.
+We focus on the entropic formulation as regularization ensures uniqueness and smoothness of the optimal coupling, which is crucial for efficient computation and gradient-based optimization [peyre2019computational].
 In the finite-sample setting, when $\mu$ and $\eta$ are sums of Dirac measures i.e.
 if $\mu=1/n\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ and $\eta=1/m\sum_{j=1}^{m}\delta_{\mathbf{y}_{j}}$, the problem above reduces to
-$\mathbf{W}_{\mathbf{C},\varepsilon}:=\min\limits_{\mathbf{\Pi}\in\Pi}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{m}\mathbf{\Pi}_{ij}\mathbf{C}_{ij}+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)\kern 5.0pt,$ (1)
+$\min\limits_{\mathbf{\Pi}\in\Pi}\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbf{\Pi}_{ij}\mathbf{C}_{ij}+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)\kern 5.0pt,$ (1)
 where $\Pi:=\big\{\mathbf{\Pi}\in\mathbb{R}_{+}^{n\times m}\,|\,\mathbf{\Pi}\mathbf{1}_{m}=1/n\mathbf{1}_{n},\,\mathbf{\Pi}^{\top}\mathbf{1}_{n}=1/m\mathbf{1}_{m}\big\}$, $\mathbf{KL}\big(\mathbf{\Pi}\big):=\sum_{ij}\mathbf{\Pi}_{ij}\log\mathbf{\Pi}_{ij}$ and $\mathbf{C}\in\mathbb{R}^{n\times m}$ with $\mathbf{C}_{ij}:=c(x_{i},y_{j})$.
 The KL divergence is strictly convex, and $\Pi$ is a nonempty and bounded set.
-Therefore, Problem (1) admits a unique minimizer, which we denote by $\mathbf{\Pi}_{\varepsilon}(\mathbf{C})$.
+Therefore, (1) admits a unique minimizer, which we denote by $\mathbf{\Pi}_{\varepsilon}(\mathbf{C})$.
 3.2 Fair Optimal Transport
@@
 The sets $\mathcal{W}$ and $\mathcal{S}$ are assumed to be finite, and are identified, respectively, with $\{1,\dots,K_{w}\}$ and $\{1,\dots,K_{s}\}$.
-We first make an assumption on the support of the measures, which is needed to leverage several recent results in statistical optimal transport [rigollet2022sample].
+To obtain theoretical results on sample complexity, we will need an assumption on the support of the measures, which is standard in statistical optimal transport [rigollet2022sample].
 {boxasm}
-The measures $\mu$ and $\eta$ are compactly supported.
-We let $p_{s}=\mathbb{P}(S=s)$ and $q_{w}=\mathbb{P}(W=w)$ the probability of each group $i$ in $\mu$ and $j$ in $\eta$, respectively, and denote by $\mathbf{p}$ and $\mathbf{q}$ the resulting distributions over $\mathcal{S}$ and $\mathcal{W}$.
+$\mu$ and $\eta$ are compactly supported.
+We denote by $p_{s}=\mathbb{P}(S=s)$ and $q_{w}=\mathbb{P}(W=w)$ the probability of each group $i$ in $\mu$ and $j$ in $\eta$, respectively, and by $\mathbf{p}$ and $\mathbf{q}$ the resulting distributions over $\mathcal{S}$ and $\mathcal{W}$.
 Fairness targets.
-Figure 1: Summary of the fairness issues arising in optimal transport when sensitive features are highly correlated with features used for cost computations.
-On the left of the figure, we show an illustrative setting in which we wish to match two populations (circles and triangles).
-Group membership is highly correlated with the individual’s location in the $(x,y)$ plane.
-This results in a block sparse transport plan, displayed on the right.
 Optimal transport imposes structure on the coupling of two probability distributions through the cost function, making it more likely to match points whose transportation cost to one another is low.
@@
 Example 3.1.
-Consider a school assignment system, in which students are characterized by their geographic position $X\in\mathbb{R}^{2}$ and a social status $S\in\{\textnormal{high},\textnormal{low}\}$.
-Similarly, schools are characterized by geographic positions $Y\in\mathbb{R}^{2}$ and a prestige level $W\in\{\textnormal{elite},\textnormal{non-elite}\}$.
-Suppose that students with $S=\textnormal{high}$ are more likely to live near elite schools, which are often situated in privileged neighborhoods, while students with low social status tend to live near non-elite schools.
-In this setting, assigning students to schools using optimal transport with Euclidean cost—minimizing the distance between students and schools—will result in highly segregated schools.
+Consider a school assignment system, in which students have a location $X\in\mathbb{R}^{2}$ and social status $S\in\{\textnormal{high},\textnormal{low}\}$, and schools have a location $Y\in\mathbb{R}^{2}$ and prestige $W\in\{\textnormal{elite},\textnormal{non-elite}\}$.
+If high-status students tend to live near elite schools and low-status students near non-elite schools, then optimal transport with Euclidean cost—minimizing the distance between students and schools—will produce highly segregated schools.
 In this example, the block-sparse structure of the optimal transport plan can be seen as a source of unfairness, as the matching will be highly correlated with the social status of the students and the elitist nature of the schools.
-To address such issues and define a fair transport plan, we are given a fairness target $\mathbf{F}$, which is a $K_{s}\times K_{w}$ matrix that specifies, for each pair of groups $(s,w)\in\mathcal{S}\times\mathcal{W}$, the desired probability of matching members of group $s$ with group $w$.
-To be valid, the matrix $\mathbf{F}$ should itself be a coupling between $\mathbf{p}$ and $\mathbf{q}$: it should be non-negative and satisfy the constraints
-$\sum\limits_{w=1}^{K_{w}}\mathbf{F}_{sw}=p_{s}\textnormal{ and }\sum\limits_{s=1}^{K_{s}}\mathbf{F}_{sw}=q_{w}$
-for all $(s,w)$. We hence write $\Pi(\mathbf{p},\mathbf{q})$ the set of fairness targets.
+To address such issues and define fair transport plans, we are given a fairness target $\mathbf{F}$, which is a $K_{s}\times K_{w}$ matrix that specifies, for each pair of groups $(s,w)\in\mathcal{S}\times\mathcal{W}$, the desired probability of matching members of group $s$ with group $w$.
+To be valid, the matrix $\mathbf{F}$ should itself be a coupling between $\mathbf{p}$ and $\mathbf{q}$: it should be non-negative and satisfy the following constraints for all $(s,w)$:
+$\textstyle\sum{w=1}^{K_{w}}\mathbf{F}_{sw}=p_{s}\textnormal{ and }\sum{s=1}^{K_{s}}\mathbf{F}_{sw}=q_{w}\kern 5.0pt.$
+We hence write $\Pi(\mathbf{p},\mathbf{q})$ the set of fairness targets.
 Example 3.2.
-Consider the segregated schooling system introduced in Example 3.1.
+Consider the segregated schooling system of Example 3.1.
 Let $p:=\mathbb{P}(S=\textnormal{low})$ and $q:=\mathbb{P}(W=\textnormal{non-elite})$.
-A newly appointed city administrator wishes to strongly limit social homogamy within public schools, and hence requires $60\%$ of student with low social status to be matched to elite schools.
+A newly appointed city administrator wishes to limit social homogamy within public schools, and hence requires $60\%$ of student with low social status to be matched to elite schools.
 The fairness target can be written as
-$\mathbf{F}=\begin{bmatrix}0.4\times p&0.6\times p\\ q-0.4\times p&1-q-0.6\times p&\end{bmatrix}$.
-Then, we define a coupling to be $\mathbf{F}$-fair if the amount of mass moved from group $s$ to group $w$ matches the fairness target $\mathbf{F}_{sw}$.
-Definition 3.3 ($\mathbf{F}$-Fair coupling).
-Let $\mu\in\mathcal{P}(\mathcal{X}\times\mathcal{S})$ and $\eta\in\mathcal{P}(\mathcal{Y}\times\mathcal{W})$. We define the $\mathbf{F}$-fair transport plan between $\mu$ and $\eta$ as
-$\pi^{\star}_{c,\varepsilon,\textnormal{fair}}\in\operatorname*{arg\,min}\limits_{\pi\in\Pi_{\textnormal{fair}}(\mu,\eta)}\int c(x,y)\,d\pi(x,y)+\varepsilon\mathbf{KL}(\pi|\mu\otimes\eta)\kern 5.0pt.$ (2)
-where the set of fair couplings $\Pi_{\textnormal{fair}}(\mu,\eta)$ is defined as
-$\Pi_{\textnormal{fair}}(\mu,\eta):=\Big\{\pi\in\Pi(\mu,\eta)\mid\forall(s,w)\in\mathcal{S}\times\mathcal{W},\,\pi_{SW}(s,w)=\mathbf{F}_{sw}\Big\}$
-where $\pi_{SW}$ is the coupling on $\mathcal{S}\times\mathcal{W}$ that is obtained from $\pi$ by marginalizing over $x$ and $y$, that is
-$\pi_{SW}(s,w):$ $=\pi(\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\})=\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi(x,u,y,t)\kern 5.0pt.$
+$\mathbf{F}=\begin{bmatrix}0.4\times p&0.6\times p\\ q-0.4\times p&1-q-0.6\times p\end{bmatrix}$.
+Fair optimal transport.
+We say that a coupling is $\mathbf{F}$-fair if the amount of mass transported from group $s$ to group $w$ equals the prescribed fairness target $\mathbf{F}_{sw}$.
+An $\mathbf{F}$-fair optimal transport plan is then an $\mathbf{F}$-fair coupling that minimizes the transport cost among all such couplings.
+Definition 3.3 ($\mathbf{F}$-Fair Optimal Transport).
+Let $\mu\in\mathcal{P}(\mathcal{X}\times\mathcal{S})$ and $\eta\in\mathcal{P}(\mathcal{Y}\times\mathcal{W})$. An $\mathbf{F}$-fair optimal transport plans between $\mu$ and $\eta$ is defined as
+$\pi^{\star}_{c,\varepsilon,\mathbf{F}}\!\in\!\!\operatorname*{arg\,min}\limits_{\pi\in\Pi_{\mathbf{F}}(\mu,\eta)}\int\!\!c(x,y)\,d\pi(x,y)\!+\!\varepsilon\mathbf{KL}(\pi|\mu\otimes\eta)$ (2)
+where the set of $\mathbf{F}$-fair couplings $\Pi_{\mathbf{F}}(\mu,\eta)$ is given by
+$\Pi_{\mathbf{F}}(\mu,\eta):=\big\{\pi\in\Pi(\mu,\eta)\mid\forall(s,w),\,\pi_{SW}(s,w)=\mathbf{F}_{sw}\big\}$
+with $\pi_{SW}$ denoting the induced coupling on $\mathcal{S}\times\mathcal{W}$ obtained from $\pi$ by marginalizing over $x$ and $y$, that is
+$\pi_{SW}(s,w):$ $=\pi(\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\})\kern 5.0pt.$
 In the finite-sample case, where we have access to two datasets $(\mathbf{x}_{i},\mathbf{s}_{i})_{i=1}^{n}\in\mathcal{X}\times\mathcal{S}$ and $(\mathbf{y}_{i},\mathbf{w}_{i})_{i=1}^{m}\in\mathcal{Y}\times\mathcal{W}$ drawn i.i.d.
-from $\mu$ and $\eta$ respectively, the fair optimal transport problem writes
-$\begin{aligned} &\min\limits_{\mathbf{\Pi}\in\Pi}\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{C}\big]+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)\\ \textnormal{s.t.}\,\,&\forall(s,w)\in\mathcal{S}\times\mathcal{W},\,\,\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{B}_{sw}\big]=\sum\limits_{i|\mathbf{s}_{i}=s}\sum\limits_{j|\mathbf{w}_{i}=w}\mathbf{\Pi}_{ij}=\mathbf{F}_{sw}\kern 5.0pt\end{aligned}$ (3)
+from $\mu$ and $\eta$ respectively, the fair optimal transport problem (2) writes
+$\begin{aligned} &\min\limits_{\mathbf{\Pi}\in\Pi}\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{C}\big]+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)\\ \textnormal{s.t.}\,&\forall(s,w),\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{B}_{sw}\big]\!=\!\!\sum\limits_{i|\mathbf{s}_{i}=s}\sum\limits_{j|\mathbf{w}_{i}=w}\mathbf{\Pi}_{ij}=\mathbf{F}_{sw}\kern 5.0pt\end{aligned}$ (3)
 where $\mathbf{B}_{sw}:=\big(\mathds{1}_{\mathbf{s}_{i}=s}\mathds{1}_{\mathbf{w}_{j}=w}\big)_{i,j}\in\{0,1\}^{n\times m}$.
-This definition may also be extended to include individual specific weights.
-The following proposition ensures that this problem is well defined.
+The following proposition ensures that the fair OT problem is well defined.
 {boxprop}
-Under Assumption 3.2, for any $\mathbf{F}\in\Pi(\mathbf{p},\mathbf{q})$, there exists a unique fair optimal transport plan.
+Under Assumption 3.2, for any target $\mathbf{F}\in\Pi(\mathbf{p},\mathbf{q})$, there exists a unique fair OT plan.
+Remark 3.1 (From probabilistic to deterministic fair transport plans).
+Given a fair entropic transport plan, a deterministic matching can be generated by sequentially iterating over the rows and, for each row, sampling a column according to the corresponding normalized distribution.
+For large enough groups, the sampled transport plan will concentrate towards a fair deterministic matching.
+Remark 3.2 (Extensions).
+Our framework naturally extends to settings in which fairness targets are imposed only on a subset of group pairs, leaving remaining matching probabilities unconstrained, as well as to cost-weighted fairness targets that, for example, balance transport costs across groups.
 3.3 Solving Fair Optimal Transport
@@
 $\mathbf{\Pi}=\textnormal{diag}\big(e^{\mathbf{f}/\varepsilon}\big)\big(\mathbf{K}\odot\mathbf{H}\big)\textnormal{diag}\big(e^{\mathbf{g}/\varepsilon}\big)\kern 5.0pt,$
-where
-$\mathbf{K}:=\big[e^{-\mathbf{C}_{ij}/\varepsilon}\big]_{ij}\textnormal{ and }\mathbf{H}:=\sum_{sw}e^{h_{sw}/\varepsilon}\mathbf{B}_{sw}\kern 5.0pt.$
+where $\mathbf{K}:=\big[e^{-\mathbf{C}_{ij}/\varepsilon}\big]_{ij}\textnormal{ and }\mathbf{H}:=\textstyle\sum_{sw}e^{h_{sw}/\varepsilon}\mathbf{B}_{sw}$.
 This result mirrors the foundational insight behind the Sinkhorn algorithm and shows that the fairness constraint can be incorporated via a simple additional projection step.
 This leads to our modified Sinkhorn algorithm, which we call FairSinkhorn, presented in Algorithm˜1 in which the function $\Phi$ is defined as
-$\Phi(\mathbf{u},\mathbf{v}):=\bigg(\textnormal{Tr}\Big[\Big(\big[\mathbf{u}_{i}\mathbf{v}_{j}\big]_{ij}\odot\mathbf{B}_{sw}\Big)\big(\mathbf{K}\odot\mathbf{B}_{sw}\big)^{\top}\Big]\bigg)_{sw}.$
-The supplementary steps with respect to the original Sinkhorn algorithm are highlighted in blue.
+$\Phi(\mathbf{u},\mathbf{v})]_{sw}:=\sum_{i\lvert s_{i}=s}\sum_{j\lvert w_{j}=w}\mathbf{u}_{i}\mathbf{v}_{j}\mathbf{K}_{i,j}$
+The additional steps relative to the original Sinkhorn algorithm are highlighted in blue.
 1:Inputs: Cost $\mathbf{C}\in\mathbb{R}^{n\times m}$, $\varepsilon>0$, fairness target $\mathbf{F}=(\mathbf{F}_{sw})_{sw}\in[0,1]^{K_{s}\times K_{w}}$, number of iterations $T$, initializations $\mathbf{u}^{(0)}\in\mathbb{R}^{n}$, $\mathbf{v}^{(0)}\in\mathbb{R}^{m}$, and $\mathbf{L}^{(0)}\in\mathbb{R}^{K_{s}\times K_{w}}$.
@@
 4:for $t=0,\dots,T-1$ do
-5:  $\mathbf{u}^{(t+1)}\leftarrow\mathbf{a}\oslash\Big((\mathbf{K}\odot\mathbf{T}^{(t)})\mathbf{v}^{(t)}\Big)$
-6:  $\mathbf{v}^{(t+1)}\leftarrow\mathbf{b}\oslash\Big((\mathbf{K}\odot\mathbf{T}^{(t)})^{\top}\mathbf{u}^{(t+1)}\Big)$
+5:  $\mathbf{u}^{(t+1)}\leftarrow n^{-1}\mathbf{1}\oslash\Big((\mathbf{K}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\odot\mathbf{T}^{(t)}})\mathbf{v}^{(t)}\Big)$
+6:  $\mathbf{v}^{(t+1)}\leftarrow m^{-1}\mathbf{1}\oslash\Big((\mathbf{K}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\odot\mathbf{T}^{(t)}})^{\top}\mathbf{u}^{(t+1)}\Big)$
 7:  $\mathbf{L}^{(t+1)}\leftarrow\mathbf{F}\oslash\Phi(\mathbf{u}^{(t+1)},\mathbf{v}^{(t+1)})$
-8:  $\mathbf{T}^{(t+1)}\leftarrow\sum_{sw}\ell^{(t+1)}_{sw}\mathbf{B}_{sw}$ where $\ell^{(t+1)}_{sw}=\big(\mathbf{L}^{(t+1)}\big)_{sw}$
+8:  $\mathbf{T}^{(t+1)}\!\leftarrow\!\sum_{sw}\ell^{(t+1)}_{sw}\mathbf{B}_{sw}$ with $\ell^{(t+1)}_{sw}\!=\!\big(\mathbf{L}^{(t+1)}\big)_{sw}$
 9:end for
@@
 4 Two Strategies for Approximately Fair OT
-While the previous section showed that exact fairness can be enforced via a projection step in FairSinkhorn, strictly satisfying fairness constraints can substantially increase transport cost, sometimes beyond what is acceptable in practice.
-This price of fairness has been studied in operations research for resource allocation problems [bertsimas2011price] and also arises in supervised learning, where enforcing exact group fairness constraints may require sacrificing predictive accuracy [zhao2019inherent].
-Motivated by these observations, it is often preferable in practice to allow some tolerance, trading off slight deviations from fairness against overall efficiency.
-In this section, we implement this idea by replacing the linear fairness constraints in (LABEL:eq:pure_empirical_fairness) with a relaxed condition
-$\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}):=\sum_{(s,w)\in\mathcal{S}\times\mathcal{W}}\bigg(\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{B}_{sw}\big]-\mathbf{F}_{sw}\bigg)^{2}\leq\rho,$ (4)
-where $\rho\geq 0$ is a tolerance level.
-This formulation allows the transport plan to deviate slightly from the exact fairness target while still controlling the total fairness violation.
-To find optimal transport plans under this relaxed constraint, we propose two strategies.
+FairSinkhorn enforces exact fairness, which can substantially increase transport cost, sometimes beyond what is acceptable in practice.
+This “price of fairness” has been studied in operations research for resource allocation problems [bertsimas2011price] and also arises in supervised learning, where enforcing exact group fairness constraints may sacrifice predictive accuracy [zhao2019inherent].
+Motivated by these observations, in this section we study approximately fair OT formulations, replacing the linear fairness constraints in (LABEL:eq:pure_empirical_fairness) with a relaxed condition that allows slight deviations from the fairness target while still controlling the total fairness violation:
+$\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}):=\sum_{(s,w)\in\mathcal{S}\times\mathcal{W}}\Big(\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{B}_{sw}\big]-\mathbf{F}_{sw}\Big)^{2}\leq\rho,$ (4)
+where $\rho\geq 0$ is a tolerance level. To find OT plans under this relaxed constraint, we propose two strategies.
 4.1 Fairness-Penalized OT
@@
 A line search then determines a convex combination of the current iterate and this direction, guaranteeing descent.
-In practice, the subproblem at iteration $t$ corresponds to an entropic OT problem with the modified cost $\mathbf{C}+\nabla\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}^{t})$ which is solved efficiently with the Sinkhorn algorithm.
+The subproblem at iteration $t$ corresponds to an entropic OT problem with the modified cost $\mathbf{C}+\nabla\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}^{t})$, which is solved efficiently with the Sinkhorn algorithm.
 This makes the algorithm computationally practical even for large-scale problems.
 Sample complexity of fairness-penalized OT.
-An important contribution of our work is to establish a sample complexity bound for the above fairness-penalized entropic optimal transport cost, by building upon results from entropic optimal transport [genevay2019sample, mena2019statistical].
-To formalize this, note that the optimization problem in (5) can be defined for arbitrary measures $\mu\in\mathcal{P}(\mathcal{X}\times\mathcal{S})$ and $\eta\in\mathcal{P}(\mathcal{Y}\times\mathcal{W})$ as (see Section˜B.2 for a formal definition)
-$m^{\star}(\mu,\eta):=\min_{\pi\in\Pi(\mu,\eta)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}c(x,y)\,d\pi(x,y)+\varepsilon\mathbf{KL}\big(\pi||\mu\otimes\eta\big)+\lambda\mathcal{L}_{\mathbf{F}}(\pi)\kern 5.0pt.$
+We establish a sample complexity bound for the above fairness-penalized optimal transport cost by building upon results from entropic optimal transport [genevay2019sample, mena2019statistical].
+To formalize this, note that the optimization problem in (5) can be defined for arbitrary measures $\mu\in\mathcal{P}(\mathcal{X}\times\mathcal{S})$ and $\eta\in\mathcal{P}(\mathcal{Y}\times\mathcal{W})$ as (see Section˜C.2 for a formal definition)
+$m^{\star}(\mu,\eta)$ $:=\min_{\pi\in\Pi(\mu,\eta)}\textstyle\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}c(x,y)\,d\pi(x,y)$
+$\qquad+\varepsilon\mathbf{KL}\big(\pi||\mu\otimes\eta\big)+\lambda\mathcal{L}_{\mathbf{F}}(\pi)\kern 5.0pt.$
 Let $\mu_{n}=\sum_{i=1}^{n}\delta_{x_{i}}$ and $\eta_{n}=\sum_{i=1}^{n}\delta_{y_{i}}$ where $(x_{i})_{1\leq i\leq n}$ and $(y_{i})_{1\leq i\leq n}$ are i.i.d.
 samples drawn from $\mu$ and $\eta$.
-In Section˜4.1, we quantify how well we can estimate $m^{\star}(\mu,\eta)$ with $m^{\star}(\mu_{n},\eta_{n})$ in function of the sample size $n$.
-Similar to [rigollet2022sample] and [genevay2019sample], our proof relies on the fact that the measures are compactly supported.
+In Section˜4.1, we quantify how well we can estimate $m^{\star}(\mu,\eta)$ with $m^{\star}(\mu_{n},\eta_{n})$ as a function of the sample size $n$.
+Similar to [rigollet2022sample] and [genevay2019sample], our proof relies on the fact that the measures are compactly supported (Assumption 3.2).
 We also require the following assumption on the ground cost function.
@@
 $\mathbb{E}\big\lvert m^{\star}(\mu_{n},\eta_{n})-m^{\star}(\mu,\eta)\big\lvert\lesssim\frac{1}{\sqrt{n}}\kern 5.0pt.$
-Compared to existing sample complexity bounds for the standard (non-penalized) entropic OT problem [genevay2019sample, mena2019statistical], our result attains the same scaling with the number of samples $n$, demonstrating that introducing the fairness penalty does not worsen the statistical efficiency of the estimator.
-The proof can be extended to generic convex penalties at the price of a supplementary $\log(n)$ factor.
+Our result matches the $O(n^{-1/2})$ scaling of sample complexity bounds for standard entropic OT [genevay2019sample, mena2019statistical], showing that adding the fairness penalty does not reduce statistical efficiency.
+Interesting, the result can be extended to generic convex penalties at the cost of a supplementary $\log(n)$ factor.
 Proof sketch for Section˜4.1..
-$Y_{n}-m^{\star}(\mu,\eta)\leq m^{\star}(\mu_{n},\eta_{n})-m^{\star}(\mu,\eta)\leq Z_{n}-m^{\star}(\mu,\eta)\kern 5.0pt.$
+Our proof works by finding two adequate random variables $Y_{n}$ and $Z_{n}$ such that
+$Y_{n}\!-\!m^{\star}(\mu,\eta)$ $\!\leq\!m^{\star}(\mu_{n},\eta_{n})\!-\!m^{\star}(\mu,\eta)\!\leq\!Z_{n}\!-\!m^{\star}(\mu,\eta).$
 Lower bound.
 We obtain $Y_{n}$ through linearization of the convex penalty, which is guaranteed to yield a proper lower bound through the connection between the optimality conditions of our penalized problem and its linearized counterpart [rakotomamonjy2015generalized].
-This linearized problem may then be cast as an entropic optimal transport problem with modified cost, allowing us to leverage standard results on the sample complexity of entropic optimal transport [genevay2019entropy, mena2019statistical].
-Upper bound. To obtain $Z_{n}$ we consider $\hat{\pi}_{n}^{\star}$, the minimizer of the linearized problem, and insert it in the fairness-penalized loss function, that is we consider
-$m^{\star}(\mu_{n},\eta_{n})-m^{\star}(\mu,\eta)\leq\langle c,\hat{\pi}_{n}^{\star}\rangle+\varepsilon\textbf{KL}(\hat{\pi}_{n}^{\star}|\mu_{n}\otimes\eta_{n})+\lambda\mathcal{L}_{F}(\hat{\pi}_{n}^{\star})-m^{\star}(\mu,\eta),$ (6)
+This linearized problem can be cast as an entropic OT problem with modified cost, allowing us to leverage existing sample complexity results [genevay2019entropy, mena2019statistical].
+Upper bound. To obtain $Z_{n}$ we evaluate the fairness-penalized loss at the minimizer $\hat{\pi}_{n}^{\star}$ of the linearized problem, that is, we consider
+$m^{\star}(\mu_{n},\eta_{n})-m^{\star}(\mu,\eta)$
+$\hskip-100.00015pt\leq\!\langle c,\hat{\pi}_{n}^{\star}\rangle\!+\!\varepsilon\textbf{KL}(\hat{\pi}_{n}^{\star}|\mu_{n}\!\otimes\!\eta_{n})\!+\!\lambda\mathcal{L}_{F}(\hat{\pi}_{n}^{\star})\!-\!m^{\star}(\mu,\eta),$ (6)
 where the upper bound follows from the fact that $\hat{\pi}_{n}^{\star}$ satisfies the constraints defining $m^{\star}(\mu_{n},\eta_{n})$.
-To establish that, in expected value, the absolute value of (6) vanishes at the rate of $1/\sqrt{n}$, we again make use of the connection between the optimality conditions of our penalized problem and its linearized counterpart [rakotomamonjy2015generalized] to reduce the problem to the sample complexity of entropic optimal transport.
-We handle the part $\lambda\mathcal{L}_{F}(\hat{\pi}_{n}^{\star})-m^{\star}(\mu,\eta)$ by considering a generalization of Theorem 6 from [rigollet2022sample].
+To establish that the expectation of the absolute value of (4.1) vanishes at the rate of $1/\sqrt{n}$, we again make use of the connection between the optimality conditions of our penalized problem and its linearized counterpart [rakotomamonjy2015generalized] to reduce the problem to the sample complexity of entropic OT.
+We handle the part $\lambda\mathcal{L}_{F}(\hat{\pi}_{n}^{\star})-m^{\star}(\mu,\eta)$ through a generalization of Theorem 6 from [rigollet2022sample].
 ∎
 4.2 Fair OT via Cost Learning
-The key idea behind our second strategy is to learn a ground cost function $c_{\theta}$ so that the entropic optimal transport (OT) solution $\mathbf{\Pi}(c_{\theta})$ induced by this cost minimizes the fairness violation.
-In other words, rather than modifying the transport plan directly, we reshape the geometry of the OT problem itself.
-To control deviations from the original OT problem, we require the learned cost to remain close to a user-chosen baseline cost $c_{\textnormal{base}}$ that encodes prior knowledge or domain-specific structure.
-This perspective naturally leads to a bilevel optimization problem, where the outer level adjusts the cost parameters and the inner level solves an OT problem whose solution depends on those parameters.
-Formally, we consider the bilevel problem
-$\min\limits_{\theta\in\Theta}\mathcal{L}_{\mathbf{F}}\big(\mathbf{\Pi}_{\varepsilon}(c_{\theta})\big)+\frac{1}{\lambda}\mathscr{D}(c_{\theta},c_{\textnormal{base}})\,\,\textnormal{ s.t.
-}\,\,\mathbf{\Pi}_{\varepsilon}(c_{\theta})=\operatorname*{arg\,min}\limits_{\mathbf{\Pi}\in\Pi}\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{C}_{\theta}\big]+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)\kern 5.0pt,$ (7)
-where $(c_{\theta})_{\theta\in\Theta}$ is a parameterized family of cost functions, $\mathbf{C}_{\theta}:=\big[c_{\theta}(\mathbf{x}_{i},\mathbf{y}_{j})\big]_{ij}$, $c_{\textnormal{base}}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}^{+}$ is a baseline cost function, and $\mathscr{D}$ is a discrepancy measure between cost functions that encourages the learned cost to remain close to the baseline.
+The key idea behind our second strategy is to learn a ground cost function $c_{\theta}$ so that the entropic OT solution $\mathbf{\Pi}(c_{\theta})$ it induces minimizes the fairness violation.
+In other words, rather than modifying the transport plan directly, we reshape the geometry of the OT problem.
+This perspective leads naturally to a bilevel optimization problem, with an outer level that adjusts the cost parameters and an inner level that solves an OT problem whose solution depends on those parameters.
+Formally, we consider
+$\begin{aligned} &\min\limits_{\theta\in\Theta}\mathcal{L}_{\mathbf{F}}\big(\mathbf{\Pi}_{\varepsilon}(c_{\theta})\big)+\frac{1}{\lambda}\mathscr{D}(c_{\theta},c_{\textnormal{base}})\\ &\textnormal{s.t.}\,\,\mathbf{\Pi}_{\varepsilon}(c_{\theta})=\operatorname*{arg\,min}\limits_{\mathbf{\Pi}\in\Pi}\textnormal{Tr}\big[\mathbf{\Pi}^{\top}\mathbf{C}_{\theta}\big]+\varepsilon\mathbf{KL}\big(\mathbf{\Pi}\big)\kern 5.0pt,\end{aligned}$ (7)
+where $(c_{\theta})_{\theta\in\Theta}$ is a parameterized family of cost functions, $\mathbf{C}_{\theta}:=\big[c_{\theta}(\mathbf{x}_{i},\mathbf{y}_{j})\big]_{ij}$, $c_{\textnormal{base}}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}^{+}$ is a baseline cost encoding prior knowledge or domain-specific structure, and $\mathscr{D}$ is a discrepancy measure that encourages the learned cost to remain close to the baseline, thereby controlling deviations from the original OT problem.
 Remark 4.1.
-This approach is related to prior work on cost learning in optimal transport [JMLR:v20:18-700], which aims to infer a cost function that explains observed matchings or couplings.
+This formulation is related to prior work on cost learning in optimal transport [JMLR:v20:18-700, andrade2023sparsistency], which aims to infer a cost function that explains observed matchings or couplings.
 Here, however, the objective is different: instead of fitting a cost to reproduce a given transport plan, we learn a cost whose induced OT solution optimizes a downstream criterion, namely fairness.
-Problem (7) is well posed because the inner problem is an entropic OT problem, which is strongly convex over the set of couplings $\Pi$ and therefore has a unique solution $\mathbf{\Pi}_{\varepsilon}(c_{\theta})$ for any choice of $\theta$.
-This uniqueness ensures that the bilevel objective is a well-defined function of the cost parameters.
-In practice, the bilevel problem can be optimized using gradient-based methods: one can differentiate through the iterative solver (iterative differentiation) [Bolte2023one_step_diff, Franceschi2017ForwardReverse, Maclaurin2015, Pauwels2023], or use (approximate) implicit differentiation to approximate the gradient of the bilevel objective [Dagreou2022SABA, Dagreou2024SRBA, Eisenberger2022, Ghadimi2018].
-Our approach can accommodate a wide range of parameterizations for $c_{\theta}$. Two concrete examples are provided below, with the Euclidean distance serving as a natural baseline in both cases.
+Problem (7) is well posed since the inner entropic OT problem is strongly convex over $\Pi$, yielding a unique solution $\mathbf{\Pi}_{\varepsilon}(c_{\theta})$ for any $\theta$ and thus a well-defined bilevel objective.
+This bilevel problem can be optimized using gradient-based methods by differentiating through an iterative solver (iterative differentiation) [Bolte2023one_step_diff, Franceschi2017ForwardReverse, Maclaurin2015, Pauwels2023] or using implicit differentiation to approximate the gradient of the bilevel objective [Dagreou2022SABA, Dagreou2024SRBA, Eisenberger2022, Ghadimi2018].
+Our approach can accommodate a wide range of parameterizations for $c_{\theta}$.
 Example 4.1 (Mahalanobis cost).
+Assume $\mathcal{X}=\mathcal{Y}=\mathbb{R}^{d}$.
 For a PSD matrix $\mathbf{M}$, the Mahalanobis cost $c_{\mathbf{M}}$ is defined by $c_{\mathbf{M}}(x,y)=(x-y)^{\top}\mathbf{M}(x-y)$.
-Intuitively, this corresponds to applying a linear transformation to $x$ and $y$ before computing the squared Euclidean distance.
+This corresponds to applying a linear transformation to $x$ and $y$ before computing the squared Euclidean distance.
 A key advantage of this parametrization is its interpretability: the learned matrix $\mathbf{M}$ directly reveals which directions or features in the data are emphasized or downweighted in the transport cost.
 Example 4.2 (Neural cost).
-A more flexible parametrization uses a neural network $\phi_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ to embed the inputs, defining the cost as $c_{\theta}(x,y)=\lVert\phi_{\theta}(x)-\phi_{\theta}(y)\rVert^{2}_{2}$.
-The network can be trained from scratch or initialized from a pretrained model and fine-tuned for the specific OT task, enabling highly expressive cost functions that capture complex relationships between $x$ and $y$.
+A more flexible parametrization uses neural networks $\phi_{\theta_{1}}:\mathcal{X}\to\mathbb{R}^{k}$ and $\phi_{\theta_{2}}:\mathcal{Y}\to\mathbb{R}^{k}$ to embed the inputs, defining the cost as $c_{\theta}(x,y)=\lVert\phi_{\theta_{1}}(x)-\phi_{\theta_{2}}(y)\rVert^{2}_{2}$.
+The networks can be trained from scratch or initialized from a pretrained model and fine-tuned for the specific OT task, enabling highly expressive cost functions that capture complex relationships between $x$ and $y$.
 Fairness on unseen samples.
@@
 This naturally raises the question of whether transport plans obtained by solving an optimal transport problem with the learned cost on previously unseen data will maintain a comparable level of fairness.
-To address this question, we provide a deviation bound on the fairness of the transport plan that holds uniformly over all cost functions in the parameterized family.
-Let $\pi_{\varepsilon}^{\star}(c_{\theta})$ denote the optimal transport plan on two measures $\mu$ and $\eta$.
+To address this question, we derive a pointwise deviation bound on the fairness of the transport plan computed from finite samples relative to the population solution, uniformly over the parameterized cost family.
+Let $\pi_{\varepsilon}^{\star}(c_{\theta})$ denote the OT plan on two measures $\mu$ and $\eta$.
 Our goal is to control the difference between the fairness of this plan, defined as
-$\mathcal{L}_{\mathbf{F}}\big(\pi_{\varepsilon}^{\star}(c_{\theta})\big):=\sum\limits_{s=1}^{K_{s}}\sum\limits_{w=1}^{K_{w}}\bigg(\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi_{\varepsilon}^{\star}(c_{\theta})(x,u,y,t)-\mathbf{F}_{sv}\bigg)^{2}\kern 5.0pt,$
+$\mathcal{L}_{\mathbf{F}}\big(\pi_{\varepsilon}^{\star}(c_{\theta})\big)$
+$\hskip-100.00015pt:=\sum\limits_{s=1}^{K_{s}}\sum\limits_{w=1}^{K_{w}}$ $\Big(\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi_{\varepsilon}^{\star}(c_{\theta})(x,u,y,t)-\mathbf{F}_{sv}\Big)^{2},$
 and the fairness $\mathcal{L}_{\mathbf{F}}\big(\mathbf{\Pi}(c_{\theta})\big)$ of the optimal transport plan $\mathbf{\Pi}(c_{\theta})$ computed on finite-sample versions $\mu_{n}$ and $\eta_{m}$.
@@
 {boxasm}
-There exists a constant $R_{\Theta}>0$ such that
-$\sup\limits_{\theta\in\Theta}\,\,\lVert c_{\theta}\rVert_{\infty}:=\sup\limits_{\theta\in\Theta}\quad\max\limits_{\begin{subarray}{c}x\in\textnormal{supp}(\mu)\\ \,y\in\textnormal{supp}(\eta)\end{subarray}}c_{\theta}(x,y)<R_{\Theta}.$
+There exists a constant $R_{\Theta}>0$ s.t.
+$\sup\limits_{\theta\in\Theta}\,\,\lVert c_{\theta}\rVert_{\infty}:=\sup\limits_{\theta\in\Theta}~~\max\limits_{x\in\textnormal{supp}(\mu),y\in\textnormal{supp}(\eta)}c_{\theta}(x,y)<R_{\Theta}.$
 This assumption is mildly restrictive.
 For instance, it holds if $\Theta$ is bounded and the function $(\theta,x,y)\mapsto c_{\theta}(x,y)$ is continuous on $\Theta\times\mathcal{X}\times\mathcal{Y}$.
-For instance, for the Mahalanobis cost function, the parameter space $\Theta$ is a subset of the space of positive definite matrices, and we have
-$R_{\Theta}=\max\limits_{\mathbf{M}\in\Theta}\lVert\mathbf{M}\rVert_{\textnormal{F}}\max\limits_{\begin{subarray}{c}x\in\textnormal{supp}(\mu)\\ \,y\in\textnormal{supp}(\eta)\end{subarray}}\lVert x-y\rVert_{2}$
-which is finite under Assumption 3.2 as long $\Theta$ is bounded. Similarly, for the neural cost model, provided that the neural network is $C_{\Theta}$-Lipschitz, we have
-$R_{\Theta}=C_{\Theta}\max\limits_{\begin{subarray}{c}x\in\textnormal{supp}(\mu)\\ \,y\in\textnormal{supp}(\eta)\end{subarray}}\lVert x-y\rVert_{2}.$
+For instance, for the Mahalanobis cost function (Example 4.1), the parameter space $\Theta$ is a subset of the space of positive definite matrices and we have
+$R_{\Theta}=\max\limits_{\mathbf{M}\in\Theta}\lVert\mathbf{M}\rVert_{\textnormal{F}}\max\limits_{x\in\textnormal{supp}(\mu),y\in\textnormal{supp}(\eta)}\lVert x-y\rVert_{2}$
+which is finite under Assumption 3.2 as long $\Theta$ is bounded. Similarly, for the neural cost function (Example 4.2), provided that the neural networks are $C_{\Theta}$-Lipschitz, we have
+$R_{\Theta}=C_{\Theta}\max\limits_{x\in\textnormal{supp}(\mu),y\in\textnormal{supp}(\eta)}\lVert x-y\rVert_{2}\kern 5.0pt.$
 {boxthm}
-[Deviation bound for fairness] Assume for simplicity that $n=m$. Under Assumptions 3.2 and 4.2, it holds
-$\sup\limits_{\theta\in\Theta}\mathbb{E}\big[\big\lvert\mathcal{L}_{\mathbf{F}}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})\big)-\mathcal{L}_{\mathbf{F}}\big(\pi^{\star}_{\epsilon}(c_{\theta})\big)\big\lvert\big]\lesssim\frac{\exp\big(5R_{\Theta}\varepsilon^{-1}\big)}{\sqrt{n}}.$
+Assume for simplicity that $n=m$. Under Assumptions 3.2 and 4.2, we have
+$\sup\limits_{\theta\in\Theta}\mathbb{E}\big[\big\lvert\mathcal{L}_{\mathbf{F}}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})\big)\!-\!\mathcal{L}_{\mathbf{F}}\big(\pi^{\star}_{\epsilon}(c_{\theta})\big)\big\lvert\big]\!\lesssim\!\frac{\exp\big(\frac{5R_{\Theta}}{\varepsilon}\big)}{\sqrt{n}}\kern 5.0pt.$
 Proof sketch for Section˜4.2.
-We generalize rigollet2022sample for any bounded cost $c_{\theta}$ by proving that for any test function $\varphi\in L^{\infty}(\mu\otimes\eta)$ and for any $t>0$, we have with probability $1-18e^{-t^{2}}$
-$\big\lvert\int\varphi d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert\lesssim\frac{\exp\big(5R_{\Theta}\epsilon^{-1}\big)\lVert\varphi\rVert_{L^{\infty}(\mu\otimes\eta)}\cdot t}{\sqrt{n}}$
+We generalize rigollet2022sample to any bounded cost $c_{\theta}$ by proving that for any test function $\varphi\in L^{\infty}(\mu\otimes\eta)$ and for any $t>0$, we have with probability $1-18e^{-t^{2}}$
+$\big\lvert\!\textstyle\int\!\varphi d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert\lesssim\displaystyle\frac{\exp\big(5R_{\Theta}\epsilon^{-1}\big)\lVert\varphi\rVert_{L^{\infty}(\mu\otimes\eta)}\cdot t}{\sqrt{n}}$
 which yield the following bound in expectation
-$\mathbb{E}\big[\big\lvert\int\varphi d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert\big]\lesssim\frac{\exp\big(5R_{\Theta}\epsilon^{-1}\big)\lVert\varphi\rVert_{L^{\infty}(\mu\otimes\eta)}}{\sqrt{n}}.$
-Then, using the identity $\lvert a^{2}-b^{2}\rvert\leq\lvert a+b\rvert\lvert a-b\rvert$, we show that
-$\lvert\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}_{\epsilon}(c_{\theta}))-\mathcal{L}_{\mathbf{F}}(\pi^{\star}_{\epsilon}(c_{\theta}))\rvert\lesssim\sum_{s=1}^{K_{s}}\sum_{w=1}^{K_{w}}\lvert\textnormal{Tr}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})^{\top}\mathbf{B}_{uv}\big)-\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi^{\star}_{\epsilon}(c_{\theta})(x,u,y,t)\big\lvert$.
-For suitable choice of $\varphi_{sw}\in L^{\infty}(\mu\otimes\eta)$, we have for any $s,w\in[K_{s}]\times[K_{w}]$
-$\lvert\textnormal{Tr}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})^{\top}\mathbf{B}_{uv}\big)-\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi^{\star}_{\epsilon}(c_{\theta})(x,u,y,t)\big\lvert=\big\lvert\int\varphi_{sw}d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert$
-Putting all together yields the result. ∎
-This bound provides a uniform control over the expected difference between the fairness loss evaluated on the population and its finite-sample approximation, for any cost function $c_{\theta}$ in the parameterized family.
+$\mathbb{E}\big[\big\lvert\textstyle\int\varphi d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert\big]\lesssim\frac{\exp(5R_{\Theta}\epsilon^{-1})\lVert\varphi\rVert_{L^{\infty}(\mu\otimes\eta)}}{\sqrt{n}}.$
+Then, using $\lvert a^{2}-b^{2}\rvert\leq\lvert a+b\rvert\lvert a-b\rvert$, we show that
+$\lvert\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}_{\epsilon}(c_{\theta}))\!-\!\mathcal{L}_{\mathbf{F}}(\pi^{\star}_{\epsilon}(c_{\theta}))\rvert\!\lesssim\!\sum_{s=1}^{K_{s}}\sum_{w=1}^{K_{w}}\big\lvert\textnormal{Tr}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})^{\top}\mathbf{B}_{uv}\big)$
+$\hskip-160.00024pt-\textstyle\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi^{\star}_{\epsilon}(c_{\theta})(x,u,y,t)\big\lvert$.
+Then, for a suitable choice of $\varphi_{sw}\in L^{\infty}(\mu\otimes\eta)$, we have for any $s,w\in[K_{s}]\times[K_{w}]$
+$\big\lvert\textnormal{Tr}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})^{\top}\mathbf{B}_{uv}\big)-\textstyle\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi^{\star}_{\epsilon}(c_{\theta})(x,u,y,t)\big\lvert$
+$\hskip-129.00012pt=\big\lvert\textstyle\int\varphi_{sw}d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert$.
+Putting everything together yields the result. ∎
+This bound provides a uniform control over the expected difference between the fairness loss evaluated on the population and its finite-sample approximation.
 The exponential dependence on $\varepsilon^{-1}$ is unavoidable in the worst case; see rigollet2022sample and altschuler2022asymptotics for related discussions.
-Importantly, Section˜4.2 implies that, for sufficiently large training and test samples, the transport plan induced by the learned cost achieves a level of fairness on the test samples similar to that observed on the training samples.
-In other words, the learned cost can be reused to match new samples with fairness guarantees.
+Importantly, Section˜4.2 implies that, for two sufficiently large samples, the transport plan computed with a given cost achieves a comparable level of fairness.
+In other words, the same cost function can be reused to match new samples with fairness guarantees.
 4.3 Comparison of Both Approaches
-Convexity.
-The penalized OT problem is convex, whereas the cost learning approach involves a bilevel optimization, which is non-convex and may get stuck in local minima.
-Flexibility and Interpretability.
-Cost learning offers greater flexibility: the baseline cost function and the discrepancy $\mathscr{D}$ can be chosen by the user.
-Moreover, selecting an appropriate parameterized family can make the learned cost interpretable.
-Reusability.
-Penalized OT requires solving a new problem for each batch of samples. In contrast, a learned cost can be reused to match new samples via a standard OT problem.
+Convexity. The penalized OT problem is convex, whereas the cost learning approach involves a bilevel optimization, which is non-convex and may get stuck in local minima.
+Flexibility.
+The cost learning approach restricts the transport map to entropic-regularized solutions with a parametric cost.
+This problem is thus structurally more constrained than the penalized formulation, which allows solutions outside the set of optimal couplings and is therefore more flexible.
+Reusability and interpretability.
+Penalized OT requires solving a new problem for each batch of samples.
+In contrast, a learned cost can be reused to match new samples via a standard OT problem.
+Moreover, when the cost belongs to an interpretable family, it can yield insight into how specific features contribute to (un)fairness.
 5 Experiments
-5.1 Approaches and Metrics
+5.1 Experimental Setup
+Data generation.
+Building on the illustrating example of segregated school systems (Examples 3.13.2), we consider two synthetic problems.
+For the first one (Fig.˜2, left), the positions of students ($X$) and schools ($Y$) are drawn according to a mixture of two Gaussian distributions such that privileged students ($S=1$) are closer to elite schools ($W=1$), and underprivileged students closer to non-elite schools.
+For the second one (Fig.˜2, right), privileged students and elite schools are drawn from a centered Gaussian distribution, while underprivileged students and non-elite schools are sampled on a centered circle of radius 2.
+In both cases, there are 10 times more schools than students, and sensitive attributes are balanced.
+Figure 2: Simulated datasets: Gaussians (left) and Circles (right).
+Fairness target.
+We aim to assign approximately $60\%$ of underprivileged students to elite schools, which corresponds to the fairness target $\mathbf{F}=\begin{bmatrix}0.20&0.30\\ 0.28&0.22\end{bmatrix}$.
 Approaches.
-We compare our two relaxed approaches: penalized OT and cost learning.
-For cost learning, we evaluate two parameterizations: a Mahalanobis cost (˜4.1) and a neural cost (˜4.2), where $\phi_{\theta}$ is a multilayer perceptron (MLP) or a ResNet trained from scratch.
-5.2 Approaches and Metrics
-Approaches.
-We compare our two relaxed approaches: penalized OT and cost learning.
-For cost learning, we evaluate two parameterizations: a Mahalanobis cost (˜4.1) and a neural cost (˜4.2), where $\phi_{\theta}$ is a multilayer perceptron (MLP) or a ResNet embedding.
+We compare our three approaches: FairSinkhorn , penalized OT and cost learning.
+For cost learning, we evaluate two parameterizations: a Mahalanobis cost (˜4.1) and a neural cost (˜4.2), where $\phi_{\theta_{1}}$ and $\phi_{\theta_{2}}$ are multilayer perceptrons (MLP).
+Implementation details are available in Appendix˜D.
+We also include vanilla entropic OT with varying $\varepsilon$ as a baseline.
 Metrics.
-Following standard practice in fairness research, we analyze the utility–fairness trade-off by visualizing performance in the two-dimensional (utility, fairness) plane.
-The Pareto front corresponds to all points that, for any given level of fairness, achieve the lowest level of bias.
-The Pareto front consists of points that, for a given level of fairness, achieve the lowest possible bias; points on this front are considered optimal, as improving one metric requires trading off the other.
-Bias is measured relative to the transport plan obtained by standard (non-fair) OT with the same entropic regularization, using two quantities: (a) the KL divergence between transport plans, and (b) the difference in total transport cost.
-In all experiments, the target $\mathbf{F}$ has all coefficients equal to zero, except for the anti-diagonal entries, which are set to $\frac{1}{KsKw}$.
-Implementation.
-Our experiments are conducted using the Python Optimal Transport library [flamary2021pot, flamary2024pot], which provides implementations of the Sinkhorn and generalized conditional gradient algorithms.
-In the cost-learning approach, we use iterative differentiation by backpropagating through the Sinkhorn iterations [Pauwels2023] to compute gradients of the bilevel objective using the PyTorch framework [Paszke2019].
-5.3 Comparing Exact and Penalized Approaches
-To illustrate the need to move from strict group fairness constraints to relaxed ones, we compare the transport cost and the level of fairness achieved for the transport plans obtained using our FairSinkhorn algorithm, introduced in Section˜3.2, the transport plan obtained using the original Sinkhorn algorithm, and the fairness-penalized optimal transport presented in Section˜4.1, for different values of the fairness penalty.
-In Fig.˜2, one observes that FairSinkhorn achieves perfect fairness, but with high transport cost.
-In contrast, Sinkhorn achieves low transport cost with high fairness loss.
-One observes that adopting the relaxed formulation allows to attain low fairness violations while achieving a better transport cost.
-This demonstrate that the penalization enables to tune the fairness-utility trade-off of interest.
-Figure 2: Comparison of FairSinkhorn , penalized optimal transport, and Sinkhorn on the Gaussian data introduced in Section˜5.4.
+Following standard practice in fairness research, we analyze the cost–fairness trade-off by visualizing performance in the two-dimensional (cost, fairness) plane.
+Cost is measured relative to the transport plan obtained by standard (non-fair) OT with the same entropic regularization, using the difference in total transport cost.
+Figure 3: a.
 FairSinkhorn achieves perfect fairness with a high transport cost while Sinkhorn achieves low transport cost with low fairness.
 The penalized OT interpolates between Sinkhorn and FairSinkhorn .
+b.
+Cost-fairness trade-off of our penalized and cost-learning approaches on both datasets for varying fairness penalties.
+c.
+Vanilla entropic OT with different values of $\varepsilon$ is included as a baseline.
+We also display the fairness targets reached at a few selected points.
+5.2 Comparing Exact and Penalized Approaches
+To illustrate the need to move from exact group fairness constraints to relaxed ones, we compare, on the Gaussian dataset, the (cost,fairness) trade-off achieved by the original Sinkhorn algorithm, our FairSinkhorn algorithm (Section˜3.2), and our fairness-penalized approach (Section˜4.1) for varying values of the fairness penalty.
+As shown in Fig.˜3 a., our penalized approach smoothly interpolates between Sinkhorn , which attains low transport cost but large fairness violations, and FairSinkhorn , which enforces perfect fairness at the expense of higher transport cost.
+This relaxed formulation thus enables fine-grained control over the fairness-cost trade-off.
 5.3 Comparing Penalized OT and Cost Learning
-Figure 3: Synthetic Gaussian data experiment.
-Left: 2D gaussians with binary sensitive attribute $S\in\{0,1\}$ and $W\in\{0,1\}$.
-Right: Utility/fairness tradeoffs.
-On the top, the utility metric is the KL divergence between the OT plan with and without the fairness constraint in function of the fairness regularization parameter.
-On the bottom, the difference between the transport cost with fair and unfair transport plan.
-The penalized OT correspond to our first relaxed approach while de Mahalanobis, MLP and ResNet ones correspond to the cost learning approach.
-Figure 4: Transport plan we get using different level of fairness parameter for each approach with the Gaussian data.
-We now evaluate both approaches on simulated datasets.
-We consider two types of data.
-First, we study mixtures of Gaussians, where each group is assigned to a distinct mixture component (see Fig.˜3, left).
-Second, we consider a dataset composed of concentric circles (see Fig.˜5, left), which allows us to examine the effect of data geometry.
-In both settings, the penalized OT approach achieves a better utility–fairness trade-off, benefiting from its convex formulation.
-Moreover, for the cost-learning approach, we observe that using more expressive cost models enables lower levels of fairness loss to be attained.
-Figure 5: Synthetic circular data experiment.
-Left: 2D gaussians with sensitive attribute $S\in\{0,1,2\}$ and $W\in\{0,1,2\}$.
-Right: Utility/fairness tradeoffs.
-On the top, the utility metric is the KL divergence between the OT plan with and without the fairness constraint in function of the fairness regularization parameter.
-On the bottom, the difference between the transport cost with fair and unfair transport plan.
-The penalized OT correspond to our first relaxed approach while de Mahalanobis, MLP and ResNet ones correspond to the cost learning approach.
-Figure 6: Transport plan we get using different level of fairness parameter for each approach with the circular data.
+Fig.˜3 b.
+compares our two relaxed approaches across our two datasets, allowing us to highlight the effect of data geometry.
+For the Gaussian dataset, the cost-fairness trade-off of penalized OT, cost learning with a Mahalanobis metric, and cost learning with an MLP are similar.
+On the Circles dataset, however, the penalized approach achieves a superior trade-off, reaching any level of fairness at lower cost than the cost learning methods.
+This is expected, as penalized OT imposes no restrictions on the transport plans (see Section 4.3).
+Moreover, cost learning with the Mahalanobis parameterization cannot reduce fairness loss below $10^{-2}$, whereas the MLP variant is flexible enough to transform the data nonlinearly, which is required to closely match the fairness target on this problem.
+Finally, for both datasets, increasing the entropic penalty $\varepsilon$ in vanilla OT fails to achieve the desired fairness, highlighting the relevance and effectiveness of our approaches.
+Figure 4: Left: Inference time of penalized OT and cost-learning approaches, highlighting the much faster inference of the cost-learning methods once the cost function is learned.
+Right: Fairness levels achieved on new samples using the learned cost function, with vanilla OT shown as a baseline.
+Leveraging the learned cost allows new samples to attain fairness comparable to that observed during training.
 5.4 Assessing the Reusability of Learned Costs
-One of the peaks of our cost learning driven approach is that once learned, the cost function can be called on new samples without any retraining.
-In contrast, the penalized optimal transport plan cannot be reused, and we need to solve a new transport problem for every new batch of data points.
-We quantify the reusability of our bilevel algorithm in the following way.
-We first learn a Mahalanobis cost function $\widehat{\mathbf{M}}$ on a set of $n_{0}=200$ datapoints with $k_{s}=k_{w}=2$ distinct classes.
-We subsequently sample $k=50$ new datapoints from the same distribution, and $a)$ solve the penalized OT problem on this fresh dataset $b)$ solve the optimal transport problem using $\widehat{\mathbf{M}}$ to define the cost and compute the fairness of the resulting transport plan.
-Results are displayed in Figure 7.
-One observes that using the learned cost allows new data points to be matched with significantly lower inference time than the penalized approach, since the latter requires solving the entire penalized optimal transport problem each time new samples are added to the dataset.
-Figure 7: We assess the recyclability of the learned cost function by $a)$ learning a cost function on $n_{0}=200$ samples and $b)$ computing the entropic transport plan on batches of $n_{k}$ new unseen samples.
-Concurrently to this setup, we train the penalized transport plan from scratch for every new batch.
-Here, $\varepsilon=2$, $n_{k}=100$, $d=2$ and $k_{s}=k_{w}=2$.
-We set the penalty for all three approaches to yield comparable levels of fairness, but emphasize that our goal in this experiment is to compare the running times of the different approaches.
+A key advantage of our cost learning approach is that, once the cost function is learned, it can be applied to match new samples using vanilla OT without retraining.
+In contrast, the penalized optimal transport plan cannot be reused, requiring a new penalized OT problem to be solved for each incoming batch of data.
+In this section, we quantify the benefits of reusing the learned cost on new samples in terms of inference time and fairness generalization.
+To this end, we first learn a Mahalanobis cost and an MLP-based cost on the Gaussian problem, using a training dataset of $1000$ samples from $X$ and $100$ samples from $Y$.
+We subsequently draw test datasets of $500$ new samples from $X$ and $50$ new samples from $Y$, solve the vanilla entropic OT problem using the learned costs, and record both the inference time and fairness loss of the resulting transport plan for each test dataset.
+For comparison, we also report the inference time of penalized OT and the fairness of vanilla OT on the same test data.
+Results are shown in Figure 4.
+We observe that using the learned cost allows matching new samples with significantly lower inference time than the penalized approach since the latter requires solving the full penalized optimal transport problem each time new samples are added to the dataset.
+Furthermore, fairness levels on new samples remain close to those achieved during training.
+While a generalization gap is present—larger for the MLP cost than for the Mahalanobis cost, as expected—this demonstrates that the learned costs generalize to new samples and achieve substantially better fairness than vanilla OT.
 6 Conclusion
-In this paper, we introduced a novel notion of group fairness for optimal transport.
-Building on this new notion, we proposed an algorithm that solves entropic regularized optimal transport with perfect fairness, though this can lead to solutions with high transport cost.
-To achieve a better utility–fairness trade-off, we then propose two relaxed approaches: (i) a convex penalization of the entropic optimal transport objective, and (ii) a cost-learning-based method.
-For the penalized approach, we derive a sample complexity bound.
-For the cost learning approach, we bound the expected deviation of the fairness loss when the transport plan is estimated from a finite number of samples drawn from the underlying distributions.
-Deriving a generalization bound for the cost learning approach is left for future work.
+We introduced a novel notion of group fairness for optimal transport and propose an algorithm that computes entropic regularized transport with perfect fairness.
+To improve the cost–fairness trade-off, we developed two relaxed approaches: (i) a convex penalization of the entropic OT objective, and (ii) a cost-learning-based method.
+Both approaches are supported by theoretical results.
+We hope our work will foster further research on fairness constraints in optimal transport.
+A promising direction for future work is to study how fairness constraints in OT can lead to fairer outcomes downstream.
 Acknowledgments.
@@
 We thank the anonymous reviewer and all participants of the Eurips workshop Unifying Perspectives on Learning Biases whose comments improved this manuscript.
-Appendix A A Modified Sinkhorn Algorithm for Fair Optimal Transport
+Appendix A Extended Related Work
+A.1 Fairness in Matching Mechanisms
+Since the pioneering work of gale1962college on the stable matching problem, a series of works have provided theoretical studies of the fairness in matching mechanisms.
+A first line of research focuses on the stability of GS-like algorithms under constraints akin to individual fairness, ensuring that similar individuals receive similar outcomes by the decision-making process.
+For instance, karni2021fairness analyze the Gale-Shapley (GS) algorithm through the lens of preference informed individual fairness.
+This criterion allows for small deviations from individual fairness, which requires that similar individuals should be treated similarly.
+They devise a fair variant of the GS algorithm that ensure both approximately fair and stable matchings.
+Similarly, devic2023fairness analyze merit-based fairness of one-to-one matching mechanisms, a notion close to individual fairness, under partial uncertainty over individual’s merit.
+A second line of research introduces metrics of fairness closer to group fairness, with a specific focus on kidney paired donation programs, a singular matching problem subject to domain specific constraints — in this setting, donors and patients must be matched by forming cycles due to legal constraints, which leads to a reformulation of the matching problem as a constrained integer programming problem.
+In their seminal work, ashlagi2014free unveil an unfair market collapse phenomenon in kidney pair donation programs: instead of introducing all patients to the matching system, hospitals internally operate easy matches i.e.
+do not participate fully in the exchange mechanism.
+This leads to a loss in transplantation opportunities for highly-sensitized patients that are hard to match.
+dickerson2014price and following work by zhang2025fairness also consider the case of kidney paired donations programs.
+Their definition of fairness enforces a minimal amount of program participation for the group of highly-sensitized patients.
+More recently, lobo2025fair consider a combinatorial multimatching problem under uncertainty: in their setup, the goal is to match agents and resources without exact knowledge of their pairwise valuations.
+Their definition of fairness relies on the average utility within predefined groups, which is closest to our definition among previous work.
+Our contribution differentiates from these works in two aspects.
+First, we do not consider iterative matching mechanisms that operate at the individual level such as the GS algorithm: our work focuses on matching through (entropic) optimal transport.
+Secondly, our definition of fairness is different from previously introduced definitions, and highly modular to accommodate fairness preferences of a central planner.
+Instead of requiring individual fairness, building on individual rankings over possible assignements or participation constraints, we consider a matching problem between two populations (f.i.
+students and universities) further structured by a sensitive attribute (f.i.
+privileged and unprivileged students, and elite and non-elite universities).
+We enforce matching quotas given by a central planner, that specify for each subgroup the proportion of individuals that should be matched to subgroups in the second population.
+A.2 Fairness in Bipartite Graph Matching
+A closely related topic to our working is bipartite matching in graphs [godsil1981matchings, zdeborova2006number, noiry2021online].
+Given a graph $\big(\mathcal{U}\cup\mathcal{V},\mathcal{E}\big)$ made of two sets of vertices $\mathcal{U}$ and $\mathcal{V}$ and edges $\mathcal{E}=(e_{ij})_{i\in\mathcal{U},j\in\mathcal{V}}$, the goal is to select a one-to-one matching between vertices in $\mathcal{U}$ and vertices in $\mathcal{V}$, under the constraint that two vertices can only be matched if they are linked by an edge in $\mathcal{E}$.
+Some works have considered fairness constraints in this setup.
+For instance, castera2024price consider this problem with ad-hoc partitions of one side of the graph in disjoint groups, and study the maximal number of individuals that may be matched per group.
+This is similar to definitions of fairness through participation ratios discussed earlier in the work of zhang2025fairness.
+Similarly, panda2024individual study a doubly constrained setup in which the set of items $\mathcal{U}$ is structured through groups.
+They enforce both individual-fairness and group-fairness inspired constraints.
+In their work, $(a)$ individuals in $\mathcal{U}$ have preferences over individuals in $\mathcal{V}$, and $b)$ every individual in $\mathcal{V}$ can only accept an certain number of individuals from each subgroup in $\mathcal{U}$.
+Bipartite matching is hence fundamentally different from our setting and somewhat closer to the literature on fair division [steinhaus1949division, weller1985fair, budish2011combinatorial] since $a)$ we allow for mass splitting between individuals on both sides ; $b)$ individuals are characterized by features in our setting, and not by their relative position within a graph $\mathcal{E}$ ; $c)$ our problem is symmetric in the sens that the two matched measures play the same role.
+A.3 Fairness and Optimal Transport
+Recent work has drawn many fruitful connection between fairness and optimal transport, primarily by using OT as a tool to obtain or characterize fair learning algorithms [gordaliza2019obtaining, gouic2020projection, chzhen2020fair, DBLP:conf/aaai/ChiappaJSPJA20, DBLP:conf/nips/BuylB22, hu2023fairness, DBLP:conf/icml/XianY023, chowdhary2024fairness, xiong2024fair, divol2024demographic].
+A central insight of this line of work is that learning an optimal fair predictor can be formulated as a Wasserstein barycenter problem [gouic2020projection, chzhen2020fair, divol2024demographic].
+This perspective is orthogonal to ours: OT serves as an auxiliary mechanism for analyzing and enforcing fairness in downstream prediction tasks, whereas our goal is to study and enforce fairness of the optimal transport itself.
+Closer in spirit to our work is the recent article by nguyen2024towards, which imposes fairness constraints directly on a Wasserstein barycenter by requiring approximate equality of sliced-Wasserstein distances SW [nadjahi2020statistical] to multiple marginals $\mu_{1},\dots,\mu_{K}$.
+$\min\limits_{\mu}\sum\limits_{k=1}^{K}\textnormal{{SW}}(\mu,\mu_{k})\textnormal{ s.t }\sum\limits_{i=1}^{K}\sum\limits_{j=1}^{K}\big\lvert\textnormal{{SW}}(\mu,\mu_{i})-\textnormal{{SW}}(\mu,\mu_{j})\big\lvert\leq\varepsilon.$
+where $\textnormal{{SW}}(\cdot,\cdot)$ is the sliced-Wasserstein distance between distributions [nadjahi2020statistical] and $\varepsilon>0$ is a tolerance threshold.
+Similarly to their work, we enforce additional constraints on the optimal transport to match it to a definition of fairness based on subgroups.
+However, we study the problem of computing transport plans that satisfy explicit mass constraints between groups defined by sensitive attributes, a setting that, to the best of our knowledge, has not been previously considered.
+A.4 Constrained and Statistical Optimal Transport
+Our work borrows from recent developments in constrained optimal transport, which seeks to enforce structural properties on the transport plan [courty2016optimal, blondel2018smooth, paty2019subspace, liu2022sparsity]; for instance, korman2015optimal analyze a variant of optimal transport in which the amount of mass that can be transported between two units is upper bounded.
+Specifically, we build on rakotomamonjy2015generalized, genevay2019sample and rigollet2022sample to derive finite sample guarantees for a penalized optimal transport problem, and to bound the deviations of fairness with learned cost functions.
+Appendix B A Modified Sinkhorn Algorithm for Fair Optimal Transport
 The following is an immediate consequence of first order conditions on the Lagrangian of Problem LABEL:eq:pure_empirical_fairness.
@@
 $\mathbf{H}:=\sum\limits_{s=1}^{K_{s}}\sum\limits_{w=1}^{K_{w}}\exp({h_{sw}/\varepsilon})\odot\mathbf{B}_{sw}.$
-A fair optimal transport plan can hence be obtained using a modified Sinkhorn algorithm which we call FairSinkhorn. Define the map
-$\Phi(\mathbf{u},\mathbf{v}):=\bigg(\textnormal{Tr}\Big[\Big(\big[\mathbf{u}_{i}\mathbf{v}_{j}\big]_{ij}\odot\mathbf{B}_{sw}\Big)\big(\mathbf{K}\odot\mathbf{B}_{sw}\big)^{\top}\Big]\bigg)_{sw}.$
-The algorithm FairSinkhorn combines projections on the set of marginal constraints with a projection on the set of fairness constraints, highlighted in blue.
-Remark 1.
-Since FairSinkhorn alternates between projections on sets defined by linear constraints, it is guaranteed to converge [benamou2015iterative].
-Appendix B Proofs
-B.1 Proof of Section˜3.2
-Lemma B.1.
+Appendix C Proofs
+C.1 Proof of Section˜3.2
+Lemma C.1.
 Let $\mu\in\mathcal{P}(\mathcal{X}\times\mathcal{S})$ and $\mathcal{\eta}\in\mathcal{P}(\mathcal{Y}\times\mathcal{W})$.
@@
     $=F(S=s,W=w).$
-Proposition B.1.
+Proposition C.1.
 Assume that $\mu$ and $\eta$ have bounded support.
@@
 which shows that $\pi\in\Lambda$.
-To finish, just note that $\Pi_{\text{fair}}^{\mathbf{F}}=\Pi(\mu,\eta)\cap\Lambda$ is the intersection of two compact sets (with respect to weak topology) and Lemma B.1 establishes that it is non-empty.
+To finish, just note that $\Pi_{\text{fair}}^{\mathbf{F}}=\Pi(\mu,\eta)\cap\Lambda$ is the intersection of two compact sets (with respect to weak topology) and Lemma C.1 establishes that it is non-empty.
 Continuity of the map defining the transport problem is enough to conclude the existence of a minimizer.
@@
 ∎
-B.2 Proof of Section˜4.1
+C.2 Proof of Section˜4.1
 Notation.
@@
 $\widehat{m}^{\star}_{\infty}:=\min_{\begin{subarray}{c}\pi_{1}=\mu\\ \pi_{2}=\eta\end{subarray}}\langle c+2\lambda\sum_{s,w\in[K_{s}]\times[K_{w}]}\langle\xi_{sw},\pi_{\infty}^{\star}\rangle\xi_{sw},\pi\rangle+\varepsilon\mathbf{KL}(\pi||\mu\otimes\eta).$
-To finish the lower bound, note that $\mathbb{E}[|Y_{n}-m^{\star}_{\infty}|]\leq\mathcal{O}(\sqrt{n}^{-1})$ follows from (15) provided we show that $m_{\infty}^{\star}=\widehat{m}_{\infty}^{\star}-\lambda\mathcal{L}_{\mathbf{F}}(\pi_{\infty}^{\star})$; this is the content of Lemma B.2 that can be found in Appendix B.2.1.
+To finish the lower bound, note that $\mathbb{E}[|Y_{n}-m^{\star}_{\infty}|]\leq\mathcal{O}(\sqrt{n}^{-1})$ follows from (15) provided we show that $m_{\infty}^{\star}=\widehat{m}_{\infty}^{\star}-\lambda\mathcal{L}_{\mathbf{F}}(\pi_{\infty}^{\star})$; this is the content of Lemma C.2 that can be found in Appendix C.2.1.
 Upper bound.
@@
 $Z_{n_{1}}=\widehat{m}_{n}^{\star}\quad\text{ and }\quad Z_{n_{2}}=\lambda\mathcal{L}_{\mathbf{F}}(\hat{\pi}_{n}^{\star})-2\lambda\sum_{(s,w)\in[K_{s}]\times[K_{w}]}\big\langle\langle\xi_{sw},\pi^{\star}_{\infty}\rangle\xi_{sw},\hat{\pi}_{n}^{\star}\big\rangle.$
-Lemma B.2 shows that $m^{\star}_{\infty}=\widehat{m}_{\infty}^{\star}-\lambda\mathcal{L}_{F}(\pi^{\star}_{\infty}\rangle$ and, similar to the lower bound part of the proof (see (15)), the sample complexity of entropic optimal transport implies that,
+Lemma C.2 shows that $m^{\star}_{\infty}=\widehat{m}_{\infty}^{\star}-\lambda\mathcal{L}_{F}(\pi^{\star}_{\infty}\rangle$ and, similar to the lower bound part of the proof (see (15)), the sample complexity of entropic optimal transport implies that,
 $\mathbb{E}\big[\big|Z_{n_{1}}-\widehat{m}_{\infty}^{\star}\big|\big]=\mathbb{E}\big[\big|\widehat{m}_{n}^{\star}-\widehat{m}_{\infty}^{\star}\big|\big]\leq\mathcal{O}(\sqrt{n}^{-1}).$
@@
 $\mathbb{E}\Big[2\lambda\big|\sum_{(s,w)\in[K_{s}]\times[K_{w}]}\big\langle\langle\xi_{sw},\pi^{\star}_{\infty}\rangle\xi_{sw},\pi_{\infty}^{\star}-\hat{\pi}_{n}^{\star}\big\rangle\big|\Big]\leq\mathcal{O}(1/\sqrt{n}).$ (18)
-To obtain these two inequalities and complete the proof, it suffices to note that the conclusion of Lemma B.2 allows us to apply Lemma˜B.4 (sub-Gaussian concentration, which is stronger, readily implies the expectation bound used here), which yields the desired inequality (17).
+To obtain these two inequalities and complete the proof, it suffices to note that the conclusion of Lemma C.2 allows us to apply Lemma˜C.4 (sub-Gaussian concentration, which is stronger, readily implies the expectation bound used here), which yields the desired inequality (17).
 To obtain inequality (18) just note that, for $s,w\in[K_{s}]\times[K_{w}]$,
@@
 $\big|\langle\xi_{sw},\pi^{\star}_{\infty}\rangle^{2}-\langle\xi_{sw},\hat{\pi}_{n}^{\star}\rangle^{2}\big|\leq 2\|\xi_{sw}\|_{\infty}\big|\langle\xi_{sw},\pi^{\star}_{\infty}-\hat{\pi}_{n}^{\star}\rangle\big|;$
-and the bound in expectation is once more a matter of invoking Lemma˜B.4. ∎
-B.2.1 Technical Lemmas
-Lemma B.2.
+and the bound in expectation is once more a matter of invoking Lemma˜C.4. ∎
+C.2.1 Technical Lemmas
+Lemma C.2.
 Let $\pi^{\star}_{\infty}$ be a minimizer of
@@
 Hence, $\pi_{\infty}^{\star}$ is also a minimizer of Eq.˜20. ∎
-B.3 Proof of Section˜4.2
+C.3 Proof of Section˜4.2
 We begin by showing the following Lemma, which mimics the arguments of rigollet2022sample.
-Lemma B.3.
+Lemma C.3.
 Let $f_{\star},g_{\star}$ be dual potentials such that $\int g_{\star}(y)d\eta(y)=0$. Then we have
@@
 Using this Lemma, we get the following result which extends Theorem 6 from rigollet2022sample to parametrized costs.
-Lemma B.4.
+Lemma C.4.
 Let $\varphi\in L^{\infty}(\mu\otimes\eta)$. For all $t>0$, with probability at least $1-18e^{-t^{2}}$, we have
@@
 $\varphi_{sw}~=~\xi_{sw}+~\mathbf{F}_{sw}$
-where $\xi_{sw}$ is defined by Equation (10). We observe that $\varphi_{sw}\in L^{\infty}(\mu\otimes\eta)$. Thus, by using Lemma˜B.4 we have with probability at least $1-18e^{-t^{2}}$
+where $\xi_{sw}$ is defined by Equation (10). We observe that $\varphi_{sw}\in L^{\infty}(\mu\otimes\eta)$. Thus, by using Lemma˜C.4 we have with probability at least $1-18e^{-t^{2}}$
 $\big\lvert\textnormal{Tr}\big(\mathbf{\Pi}_{\epsilon}^{\top}(c_{\theta})\mathbf{B}_{sw}\big)-\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi^{\star}_{\epsilon}(c_{\theta})(x,u,y,t)\big\lvert$ $=\big\lvert\int\varphi_{sw}d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\lvert$
@@
 As a consequence
-$\sup_{\theta\in\Theta}\mathbb{E}[\lvert\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\mathcal{L}_{\mathbf{F}}(\pi^{*}_{\epsilon}(c_{\theta}))\rvert]\lesssim\frac{\exp(5R_{\Theta})\epsilon^{-1}}{\sqrt{n}}$
-Appendix C Computational Details
-C.1 Implementation Details
-C.1.1 Fair Sinkhorn
-We implement the FairSinkhorn algorithm from scratch using Sinkhorn-Knopp matrix scaling as proposed in cuturi2013sinkhorn.
-C.1.2 Penalized Optimal Transport
-We leverage the Python OT (POT) library [flamary2021pot] throughout our experiments.
+$\sup_{\theta\in\Theta}\mathbb{E}[\lvert\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\mathcal{L}_{\mathbf{F}}(\pi^{*}_{\epsilon}(c_{\theta}))\rvert]\lesssim\frac{\exp(5R_{\Theta})\epsilon^{-1}}{\sqrt{n}}.$
+Appendix D Computational Details
+D.1 Implementation Details
+D.1.1 Fair Sinkhorn
+We implement the FairSinkhorn algorithm by adapting the Sinkhorn-Knopp [Sinkhorn1967] implementation of the Python OT (POT) library [flamary2021pot, flamary2024pot].
+D.1.2 Penalized Optimal Transport
+We leverage the POT library [flamary2021pot, flamary2024pot] throughout our experiments.
 To solve the penalized problem, we leverage the generic conditional gradient algorithm implemented in this package.
@@
 G0 Initialized via a solution of the entropic OT problem
-numIterMax $10$
+numIterMax $2000$
 numInnerItermax $200$
@@
 The algorithm is used in conjunction with line_search_armijo provided in Python OT.
-C.1.3 Cost Learning
-We implement bilevel cost learning in Pytorch.
+D.1.3 Cost Learning
+We implement bilevel cost learning in Pytorch [Paszke2019].
 The inner optimal transport problem is solved using the ot.sinkhorn method from POT, with a warm-started initialization of the dual variables, a stoping threshold of $10^{-6}$ and a maximum of $1000$ iterations.
@@
 We use its default parameters, except for the learning rate.
-In all experiments, we set the penalty $\mathscr{D}(\cdot,\cdot)$ to be the squared Frobenius norm between the original and parametrized cost matrix, divided by the maximum entry of the euclidean cost matrix.
-C.2 Experimental Details
-To generate results in Figure 3, we generate $n_{x}=n_{y}=20$ samples for each subgroup of the $2$ subgroups in each distribution.
-For cost learning, we use a penalty grid of $\texttt{logspace}(-3,2,30)$ and set $\varepsilon=2$.
-The learning rate is set to $0.1$ for the Mahalanobis and ResNet cost models, and to $0.05$ for the MLP cost model.
-The MLP model has $3$ hidden layers and a output dimension of $8$.
-The dimension of the intermediate states is set to $4$.
-The same architectural parameters are used for the ResNet model.
-We use a SiLU activation for both models.
-All cost learning models are trained for maximum $300$ iterations of the outer problem.
-For the penalized OT results, we use a penalty grid of $\texttt{logspace}(0,3,20)$ and set $\varepsilon=2$.
-To generate results in Figure 5, we generate $n_{x}=n_{y}=20$ samples for each subgroup of the $3$ subgroups in each distribution.
-For cost learning, we use a penalty grid of $\texttt{logspace}(-3,2,30)$ and set $\varepsilon=2$.
-The learning rate is set to $0.1$ for all cost learning models.
-The MLP model has $3$ hidden layers and a output dimension of $8$.
-The dimension of the intermediate states is set to $16$.
-The same architectural parameters are used for the ResNet model.
-We use a SiLU activation for both models.
-All cost learning models are trained for maximum $300$ iterations of the outer problem.
-For the penalized OT results, we use a penalty grid of $\texttt{logspace}(-3,3,20)$ and set $\varepsilon=2$.
-To generate results in Figure 7, we set a fixed penalty of $0.1$, a fixed entropic penalty of $\varepsilon=2$ and a fixed learning rate of $0.01$.
-The Mahalanobis and ResNet cost model are used with identical parameters to Figure 3.
-We learn a cost model on the same dataset than for Figure 3 but with $1000$ samples per subgroup.
-We then evaluate the fairness of each cost model on small subsamples of $50$ samples per subgroup.
+In all experiments, we set the penalty $\mathscr{D}(\cdot,\cdot)$ to be the squared Frobenius norm between the original and parametrized cost matrix.
+The Mahalanobis matrix is initialized to the identity, while the MLP is pretrained with fairness penalty $\lambda=0$ to initialize the model close to the reference cost.
+D.2 Experimental Details
+In all our experiments involving MLP cost, the MLP model has $2$ hidden layers, an output dimension of $2$ and uses ReLU activation.
+Moreover, unless otherwise specified, the entropic regularization parameter $\varepsilon$ is set to 1.
+To generate results in the top part of Figure 3 b., we generate 250 samples from the group $X$ and 25 from the group $Y$.
+Half of the sample has sensitive attribute 0, the other half has sensitive attribute $1$.
+For cost learning with Mahalanobis cost, we use a penalty grid of $\texttt{logspace}(0,4,80)$ and the learning rate to $0.1$.
+For cost learning with MLP cost, we use a penalty grid of $\texttt{logspace}(0,4,80)$ and the learning rate to $0.05$.
+For the penalized OT results, we use a penalty grid of $\texttt{logspace}(0,3,80)$.
+For the vanilla OT results, we use an entropic grid of $\texttt{logspace}(0,2,20)$.
+To generate results in the bottom part of Figure 3 b., we generate 250 samples from the group $X$ and 25 from the group $Y$.
+Half of the sample has sensitive attribute 0, the other half has sensitive attribute $1$.
+For cost learning with Mahalanobis cost, we use a penalty grid of $\texttt{logspace}(1,3,80)$ and the learning rate to $0.05$.
+For cost learning with MLP cost, we use a penalty grid of $\texttt{logspace}(0,4,80)$ and the learning rate to $0.01$.
+For the penalized OT results, we use a penalty grid of $\texttt{logspace}(0,3,80)$.
+For the vanilla OT results, we use an entropic grid of $\texttt{logspace}(0,2,20)$.
+To generate results in Figure 4, we set a fixed entropic penalty of $\varepsilon=1$.
+For the penalized approach, the fairness penalty is set to 90.
+The the cost learning approach with Mahalanobis, the fairness penalty is 1000 and the learning rate is 0.1 while for the MLP, the fairness penalty is 500 and the learning rate is 0.05.
+The cost are learned using the mixture of Gaussian dataset with 1000 samples from $x$ and 100 samples from $y$.
+We then evaluate the fairness of each cost model on subsamples of $500$ samples from $x$ and $500$ samples from $y$.
 The results are computed for $10$ runs.
-For penalized OT, we tune the penalty value to $15$ to obtain similar levels of fairness, and use the same entropic penalty $\varepsilon=2$.